### v1

# Causal Inference with Missing Exposures, Missing Outcomes, and Dependence

Abstract
Missing data are ubiquitous in public health research. The missing-completely-at-random (MCAR) assumption is often unrealistic and can lead to meaningful bias when violated. The missing-at-random (MAR) assumption tends to be more reasonable, but guidance on conducting causal analyses under MAR is limited when there is missingness on multiple variables. We present a series of causal graphs and identification results to demonstrate the handling of missing exposures and outcomes in observational studies. For estimation and inference, we highlight the use of targeted minimum loss-based estimation (TMLE) with Super Learner to flexibly and robustly address confounding, missing data, and dependence. Our work is motivated by SEARCH-TB’s investigation of the effect of alcohol consumption on the risk of incident tuberculosis (TB) infection in rural Uganda. This study posed notable challenges due to confounding, missingness on the exposure (alcohol use), missingness on the baseline outcome (defining who was at risk of TB), missingness on the outcome at follow-up (capturing who acquired TB), and clustering within households. Application to real data from SEARCH-TB highlighted the real-world consequences of the discussed methods. Estimates from TMLE suggested that alcohol use was associated with a 49% increase in the relative risk (RR) of incident TB infection (RR=1.49, 95%CI: 1.39–1.59). These estimates were notably larger and more precise than estimates from inverse probability weighting (RR=1.13, 95%CI: 1.00–1.27) and unadjusted, complete case analyses (RR=1.18, 95%CI: 0.89–1.57). Our work demonstrates the utility of causal models for describing the missing data mechanism and TMLE for flexible inference.
Keywords: Causal Inference, Missing Data, Targeted Minimum Loss-based Estimation (TMLE), Ensemble Machine Learning, Super Learner, Tuberculosis, Missing at Random
1 Introduction
Missing data affect the integrity of analyses across the spectrum of public health research, including surveillance studies to estimate disease prevalence and randomized trials to establish efficacy of new medical products [1, 2, 3, 4, 5, 6, 7, 8]. A common approach is to simply exclude observations with missing data on the relevant variables. This “complete case” analysis can provide unbiased estimates if the data are missing as a result of a completely random process (i.e., missing-completely-at-random [MCAR]) [9]. However, this assumption is often unrealistic in health settings, leading to biased estimates as well as reduced statistical power (e.g., [10]). The Missing-at-Random (MAR) assumption allows missingness to depend on observed study variables and is often more plausible [9].
Suppose, for example, we aim to estimate the prevalence of a disease in the target population. On $N$ randomly sampled participants, we measure baseline covariates (e.g., age, gender, socio-economic status), but do not fully ascertain the outcome of interest. To recover the population-level outcome prevalence under MAR, we would need that the covariates capture all the common causes of measurement and outcomes as well as a positive probability of measurement within all possible covariate values. Then using standardization (a.k.a., G-computation), inverse weighting, multiple imputation, or more robust methods, we could obtain a point estimate and 95% confidence intervals using data on all $N$ participants. Instead, if we were willing to consider the stronger MCAR assumption, corresponding to no common causes of measurement and the outcome, then we could estimate disease prevalence with an empirical proportion among those measured.
This prevalence example illustrates the consequences of causal assumptions on missing data: stronger assumptions allow for simpler estimation approaches, while weaker assumptions require more complex estimation approaches. An analogous scenario arises in analyses of observational data, which are subject to confounding. In the following, we present a comprehensive framework to address common real-world challenges: confounding, missingness on multiple variables, and dependence between participants in our study. Our work builds on a long history of methods to address missing data (e.g., [11, 9, 12, 13, 14, 15]). As detailed below, we give special consideration to missing exposures. To the best of our knowledge, methods to handle missing exposure have largely focused on case-control designs, where exposure information is retrospectively collected on a sample of participants based on their outcome status (e.g., [16, 17, 4, 18, 19, 20]). Here, we focus on standard cross-sectional or longitudinal studies with missing exposures.
We also highlight the consequences of missingness on the baseline outcome when it is crucial to defining the population of interest. Suppose, for example, we are interested in studying the incidence of some disease. Our target population would be persons who are at risk of the developing the outcome and are, thereby, disease-free at baseline. In this setting, our incidence estimates would be subject to bias if there is differential measurement of outcomes at baseline. Using Counterfactual Strata Effects [21, 5, 22, 23, 24, 25, 26], we provide a framework for explicitly defining, identifying, and estimating parameters in such scenarios. Finally, we provide extensions for settings where participants are not independent and, instead, clustered within households or communities. Altogether, our presentation covers multi-source missing data, confounding, and dependence. We build-up from simple to complex examples in hope that our structured presentation is relevant a wide range of readers. We illustrate the practical relevance with the SEARCH-TB study.
2 Motivating example
SEARCH was a cluster randomized trial to evaluate a community-based approach to a Universal HIV Test-and-Treat intervention, as compared to the standard-of-care, in rural Kenya and Uganda (2013-2017; NCT01864603) [27]. Following a rapid census, all communities were offered multi-disease testing through community health campaigns with home-based follow-up for non-participants [28]. Through this mechanism, we measured demographic data (e.g., age, sex, education, and mobility), self-reported alcohol use (our primary exposure of interest), and tested for HIV infection. Due to high costs and complex logistics, evaluation of incident tuberculosis (TB) infection, a proxy for population-level transmission, was limited to SEARCH-TB, a sub-study in 9 eastern Ugandan communities [29, 30]. This sub-study was intentionally enriched for persons with HIV. Specifically, in each community, we sampled 100 households with at least one adult (15+ years) with HIV and 100 households without an adult with HIV. At baseline of the sub-study, tuberculin skin tests were administered to residents of the sampled households. One year later, follow-up tests were administered to participants who tested negative at baseline. Here, we provide an in-depth exploration of the methods to evaluate the effect of alcohol use on incident TB infection in SEARCH-TB [31]: (1) missingness on the exposure of interest, (2) missingness on the baseline outcome, crucial to defining the target population of interest, (3) missingness in the final outcome, and (4) confounding. Additionally, given our focus on infectious disease outcomes, the independent and identically distributed (i.i.d.) assumption was likely violated, and we accounted dependence among study participants.
3 Related Causal Problems: Building Complexity
Many studies feature only a subset of the challenges described above. We, thus, provide causal models and identification results for a series of hypothetical studies with increasing complexity in the hope of providing a useful reference for a broad range of real-world studies. For simplicity, we focus on defining and identifying causal parameters under a single level of the exposure, but our results naturally generalize to causal effects defined in terms of contrasts of counterfactual outcome distributions under two levels of the exposure (i.e., the average treatment effect or the causal risk ratio).
3.1 Classic point-treatment problem
First, we consider the classic “point-treatment” problem, where we have measured confounding by baseline covariates $L$, a binary exposure $A$ occurring at single time-point, and an outcome $Y$ occurring at the study’s close. This could represent a study of the effect of alcohol use ($A$) on incident TB infection ($Y$) among a representative cohort of persons without TB at baseline. The directed acylic graph (DAG) and non-parametric structural equation model (NPSEM) for such a study are given in Figure 1 and reflect the simplifying assumption that participants are independent (e.g., cannot transmit TB to one another due to geographical distance).
Under interventions on the causal model, we generate counterfactual outcomes corresponding to the research question of interest. Specifically, let $Y^{a}$ be the counterfactual outcome for a given participant if, possibly contrary-to-fact, they had exposure-level $A=a$. Then our causal target parameter $\mathbb{E}(Y^{a})$ is the counterfactual mean outcome if all study participants had exposure-level $A=a$. In our running example, $\mathbb{E}(Y^{a})=\mathbb{P}(Y^{a}=1)$ is the counterfactual risk of incident TB infection with alcohol use $A=a$.
Figure 1: Directed acyclic graph (DAG) for a classic point-treatment problem with complete measurement of the baseline covariates $L$, the exposure $A$, and the outcome $Y$. The corresponding non-parametric structural equation model (NPSEM) is given by $L=f_{L}(U_{L})$; $A=f_{A}(L,U_{A})$; $Y=f_{Y}(L,A,U_{Y})$ where $(U_{L},U_{A},U_{Y})$ represent unmeasured factors determining the values of the covariates, exposure, and outcome, respectively. On the DAG, $U^{*}$ represents unmeasured common causes of at least two variables in $(L,A,Y)$.
To identify our causal target parameter and express it as function of the distribution of the observed data $O=(L,A,Y)$, we would need there to be no unmeasured confounding, which corresponds to the assumption that the baseline covariates $L$ capture all the joint causes of the exposure $A$ and outcome $Y$ and can be represented as $Y^{a}\perp A\mid L$. Additionally, we need there to be a non-zero probability of having the exposure in all possible values of $L$: $\mathbb{P}(A=a\mid L)>0$ a.e.. Under these two assumptions, our causal target is equal to the G-computation formula: $\mathbb{E}[\mathbb{E}(Y\mid A=a,L)]$ [12]. Even if these assumptions are not reasonable (e.g., there are unmeasured confounders in Figure 1), we still have well-defined statistical estimand, on which we can focus our estimation efforts.
3.2 Missing exposures
Most real-world studies, however, depart from this idealized point-treatment problem. The first departure we consider is missingness on the exposure of interest. Continuing our running example, suppose that among our representative cohort of persons without TB at baseline, some participants did not answer questions about their alcohol use. Let $\Delta_{A}$ be an indicator that a participant has their exposure measured. If $\Delta_{A}=1$ for a participant, we observe their exposure $A$ as usual. However, if $\Delta_{A}=0$ for a participant, their exposure $A$ is not observed (i.e., equal to NA). The DAG and NPSEM for such a study are given in Figure 2. To define causal effects when the exposure is subject to missingness, we now consider counterfactuals indexed by both the exposure and its measurement indicator. Specifically, let $Y^{*}=Y^{\Delta_{A}=1,A=a}$ be the counterfactual outcome for a given participant if, possibly contrary-to-fact, their exposure were measured and was at level $A=a$.
Figure 2: DAG for a point-treatment problem with missingness on the exposure: $L$=baseline covariates, $\Delta_{A}$=indicator of exposure measurment, $A$=exposure, $Y$=outcome, and $U^{*}$=unmeasured common causes of at least two variables in $(L,\Delta_{A},A,Y)$. The corresponding NPSEM is $L=f_{L}(U_{L})$; $\Delta_{A}=f_{\Delta_{A}}(L,U_{\Delta_{A}})$, $A=f_{A}(L,\Delta_{A},U_{A})$, $Y=f_{Y}(L,\Delta_{A},A,U_{Y})$.
Then our causal target parameter $\mathbb{E}(Y^{*})$ is the counterfactual mean outcome if all participants had their exposure measured $\Delta_{A}=1$ and it were at level $A=a$. To identify this causal parameter and express it as a function of the distribution of the observed data $O=(L,\Delta_{A},A,Y)$, we would need $L$ to be sufficient to control for confounding and differential measurement. In more detail, the following assumptions are required:
  • Within values of the covariates $L$, outcomes among those with measured exposures are representative of outcomes among those without measured exposures: $Y^{*}\perp\Delta_{A}\mid L$.
  • There is a non-zero probability of exposure measurement within all possible values of $L$: $\mathbb{P}(\Delta_{A}=1\mid L)>0$ a.e..
  • Among those with measured exposures ($\Delta_{A}=1$),
    • $L$ captures all the common causes of the exposure and outcome: $Y^{*}\perp A\mid\Delta_{A}=1,L$.
    • there is a non-zero probability of being exposed to level $A=a$ within all possible values of $L$: $\mathbb{P}(A=a\mid\Delta_{A}=1,L)>0$ a.e..
If these assumptions hold, we can rewrite $\mathbb{E}(Y^{*})$ as the statistical estimand $\mathbb{E}[\mathbb{E}(Y\mid A=a,\Delta_{A}=1,L)]$ with proof in eAppendix A. As before, even if these assumptions do not hold (e.g., there are unmeasured common causes of measurement and the outcome in Figure 2), we still have a well-defined statistical estimand for estimation and inference.
3.3 Missing Exposures and Outcomes
We now add another common complication: missing outcomes. Continuing our running example, suppose that among our cohort of persons at risk of TB, some participants did not answer questions about their alcohol use and, despite best efforts, some participants could not be found at the end of the study for outcome ascertainment. To reflect this data generating process, we introduce new notation to reflect the longitudinal data setting. Let $L_{0}$ be baseline covariates, $L_{1}$ be additional covariates collected after the exposure but before outcome ascertainment, and $\Delta_{Y}$ be an indicator of outcome measurement. Specifically, if $\Delta_{Y}=1$ for a participant, we observe their outcome $Y$ as before. However, if $\Delta_{Y}=0$ for a participant, their outcome $Y$ is not observed (i.e., equal to NA). The simplified DAG and full NPSEM for such a study are given in Figure 3.
Figure 3: DAG with missingness on the exposure and outcome: $L_{0}$=baseline covariates, $\Delta_{A}$=indicator of measured exposure, $A$=exposure, $L_{1}$=time-varying covariates, $\Delta_{Y}$=indicator of measured outcome, and $Y$=outcome. For simplicity, we have omitted the $U^{*}$ node, representing unmeasured common causes of at least two variables in $(L_{0},\Delta_{A},A,L_{1},\Delta_{Y},Y)$. The NPSEM is given by $L_{0}=f_{L_{0}}(U_{L_{0}})$; $\Delta_{A}=f_{\Delta_{A}}(L_{0},U_{\Delta_{A}})$; $A=f_{A}(L_{0},\Delta_{A},U_{A})$; $L_{1}=f_{L_{1}}(L_{0},\Delta_{A},A,U_{L_{1}})$; $\Delta_{Y}=f_{\Delta_{Y}}(L_{0},\Delta_{A},A,L_{1},U_{\Delta_{Y}})$; $Y=f_{Y}(L_{0},\Delta_{A},A,L_{1},\Delta_{Y},U_{Y})$.
To define the causal effect when the exposure and outcome are subject to missingness, we now consider counterfactuals indexed by the exposure and two measurement indicators. Specifically, let $Y^{*}=Y^{\Delta_{A}=1,A=a,\Delta_{Y}=1}$ denote the counterfactual outcome under hypothetical interventions to ensure exposure measurement, “set” the exposure level to $A=a$, and ensure outcome measurement. To identify the counterfactual mean outcome $\mathbb{E}(Y^{*})$ and express it as a function of the distribution of the observed data $O=(L_{0},\Delta_{A},A,L_{1},\Delta_{Y},Y)$, we now need to account for the post-baseline covariates $L_{1}$, which act as time-dependent confounders. Specifically, $L_{1}$ are mediators of the exposure-outcome relationship, while “confounding” the measurement-outcome relationship. Therefore, we rely on sequential randomization and find a set of covariates that satisfies the backdoor criteria for each “intervention” node given the observed past [12]. As before, we need that the baseline covariates $L_{0}$ are sufficient to control for missing exposures and for confounding. In other words, we need the analogous identification assumptions given in the prior subsection. Additionally, we need that among participants with measured exposures at the level of interest (i.e., $\Delta_{A}=1$ and $A=a$),
  • the baseline and time-varying covariates ($L_{0},L_{1})$ capture all the common causes of outcomes and their measurement: $Y^{*}\perp\Delta_{Y}\mid L_{1},A=a,\Delta_{A}=1,L_{0}$.
  • there is a positive probability of outcome measurement within all possible values of the baseline and time-varying covariates: $\mathbb{P}(\Delta_{Y}=1\mid L_{1},A=a,\Delta_{A}=1,L_{0})>0$ a.e..
If these assumptions hold, we can rewrite $\mathbb{E}(Y^{*})$ in terms of the longitudinal G-computation formula: $\mathbb{E}\{\mathbb{E}[\mathbb{E}(Y\mid\Delta_{Y}=1,L_{1},A=a,\Delta_{A}=1,L_{0})\mid A=a,\Delta_{A}=1,L_{0})]\}$, shown in terms of iterated expectations and with proof in eAppendix B [12, 32, 33]. As before, even if these identification assumptions do not hold, we still have a well-defined statistical estimand to focus our estimation efforts.
3.4 Missing Exposures and Outcomes at Baseline and Follow-up
We now consider missingness on the outcome at baseline. Continuing our running example, we are interested in the effect of alcohol use on incident TB infection, but did not reach 100% of study participants for baseline TB testing. In other words, our cohort of participants without TB at baseline is subject to selection bias. To reflect this data generating process, we update our notation to have multiple outcome measures. Let $\Delta_{Y_{0}}$ be an indicator of outcome measurement at baseline, $Y_{0}$ be an indicator having the outcome at baseline, $\Delta_{Y_{1}}$ be an indicator of outcome measurement at follow-up, and $Y_{1}$ be an indicator having the outcome at follow-up. As previously introduced, the value of the baseline outcome defines the target population; specifically, we are interested in occurrence of the outcome ($Y_{1}=1$) among those who are at risk at baseline ($Y_{0}=0$). The corresponding causal models can be found in Figure 4.
Figure 4: DAG with missingness on the exposure, the baseline outcome, and the follow-up outcome: $L_{0}$=baseline covariates, $\Delta_{A}$=indicator of measured exposure, A=exposure, $\Delta_{Y_{0}}$=indicator of measured baseline outcome, $Y_{0}$=baseline outcome, $L_{1}$=time-dependent covariates, $\Delta_{Y_{1}}$=indicator of measured outcome at follow-up, and $Y_{1}$= outcome at follow-up. For simplicity, we have omitted the $U^{*}$ node, representing unmeasured common causes of at least 2 variables in $(L_{0},\Delta_{A},A,\Delta_{Y_{0}},Y_{0},L_{1},\Delta_{Y_{1}},Y_{1})$. The NPSEM is given by $L_{0}=f_{L_{0}}(U_{L_{0}})$; $\Delta_{A}=f_{\Delta_{A}}(L_{0},U_{\Delta_{A}})$; $A=f_{A}(L_{0},\Delta_{A},U_{A})$; $\Delta_{Y_{0}}=f_{\Delta_{Y_{0}}}(L_{0},\Delta_{A},A,U_{\Delta_{Y_{0}}})$; $Y_{0}=f_{Y_{0}}(L_{0},\Delta_{A},A,\Delta_{Y_{0}},U_{Y_{0}})$; $L_{1}=f_{L_{1}}(L_{0},\Delta_{A},A,\Delta_{Y_{0}},Y_{0},U_{L_{1}})$; $\Delta_{Y_{1}}=f_{\Delta_{Y_{1}}}(L_{0},\Delta_{A},A,\Delta_{Y_{0}},L_{1},Y_{0},U_{\Delta_{Y_{1}}})$; $Y_{1}=f_{Y_{1}}(L_{0},\Delta_{A},A,\Delta_{Y_{0}},L_{1},Y_{0},\Delta_{Y_{1}},U_{Y_{1}})$.
To define the causal target parameter in this setting, we first consider the counterfactual outcome at baseline under hypothetical interventions to ensure exposure measurement, “set” the exposure level to $A=a$, and ensure outcome measurement at baseline: $Y_{0}^{*}=Y_{0}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1}$. Additionally, we consider the counterfactual outcome at follow-up under the prior interventions as well as a hypothetical intervention to ensure outcome measurement at follow-up among those at risk at baseline: if $Y_{0}=0$, set $\Delta_{Y_{1}}=1$; else set $\Delta_{Y_{1}}=0$. Thereby, this final intervention is a dynamic or personalized one (e.g., [21, 34, 35, 36]). Denote the resulting counterfactual outcome as $Y_{1}^{*}=Y_{1}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1,\Delta_{Y_{1}}=1}$.
Now we can precisely define the causal parameter in terms of the following conditional probability, which captures the counterfactual incidence of the outcome among those at risk at baseline: $\mathbb{P}(Y_{1}^{*}=1\mid Y_{0}^{*}=0)$. Due to conditioning on a counterfactual variable, such parameters are sometimes called “Counterfactual Strata Effects” and have been used in several real-data analyses [21, 5, 22, 23, 24, 25, 26]. These effects are different from Principal Strata Effects, which are defined within subgroups of latent classes that are fundamentally not observable [37, 38, 39, 40]. For example, in SEARCH-TB, principal stratification could be applied to define the effect of alcohol use on incident TB infection among the subset of participants who would have always tested regardless of their alcohol use. Instead, our interest is the effect of alcohol use on incident TB among the entire population of persons at risk.
To identify this effect, we re-express the conditional probability as
$\mathbb{P}(Y_{1}^{*}=1\mid Y_{0}^{*}=0)=\frac{\mathbb{P}(Y_{1}^{*}=1,Y_{0}^{*}=0)}{\mathbb{P}(Y_{0}^{*}=0)}$ (1)
Then given the observed data $O=(L_{0},\Delta_{A},A,\Delta_{Y_{0}},Y_{0},L_{1},\Delta_{Y_{1}},Y_{1})$, we can identify denominator and numerator, in turn. The denominator $\mathbb{P}(Y_{0}^{*}=0)$ represents the counterfactual prevalence of not having the outcome at baseline and, thus, being at risk. The causal structure for this parameter is analogous to that of Section 3.2, but with an additional measurement indicator for the outcome. Therefore, under analogous assumptions, we can identify $\mathbb{E}(Y_{0}^{*})=\mathbb{E}\big{[}\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\big{]}$ with proof in eAppendix C.1. Since we are interested the counterfactual probability of being at risk at baseline $\mathbb{P}(Y_{0}^{*}=0)$, our statistical estimand for the denominator becomes $1-\mathbb{E}\big{[}\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\big{]}$.
In our final causal parameter (Eq. 1), the numerator $\mathbb{P}(Y_{1}^{*}=1,Y_{0}^{*}=0)$ represents the counterfactual probability of having the outcome at follow-up but not at baseline. For ease of notation, let $Z^{*}=\mathbb{I}(Y_{1}^{*}=1,Y_{0}^{*}=0)$ represent the joint indicator of these two counterfactual values. To identify $\mathbb{E}(Z^{*})=\mathbb{P}(Z^{*}=1)$, we need analogous assumptions as for the denominator together with the following. Among those known to be at risk at baseline ($\Delta_{Y_{0}}=1,Y_{0}=0$) and with measured exposure of interest ($\Delta_{A}=1,A=a$):
  • the baseline and time-varying covariates capture the common causes of the joint outcome and follow-up measurement: $Z^{*}\perp\Delta_{Y_{1}}\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}$.
  • there is a positive probability of follow-up measurement within all possible values of $L_{0}$ and $L_{1}$: $P(\Delta_{Y_{1}}=1\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})>0$ a.e..
Under these assumptions and with proof given in eAppendix C.2, the numerator is identified as
$\mathbb{P}(Z^{*}=1)\ =\ \mathbb{E}[\mathbb{E}\left(\mathbb{E}(Y_{1}\mid\Delta_{Y_{1}}=1,L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}\right)]$
Putting it all together, the statistical estimand with missing exposures, missing outcomes at baseline, and missing outcomes at follow-up is given
$\Psi(\mathbb{P};a)=\frac{\mathbb{E}[\mathbb{E}\left(\mathbb{E}(Y_{1}\mid\Delta_{Y_{1}}=1,L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}\right)]}{1-\mathbb{E}\left[\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\right]}$ (2)
for exposure level $A=a$. Then we can define associations in terms of contrasts $\Psi(\mathbb{P};a)$ at different exposure levels. Concretely, in SEARCH-TB, we were interested in evaluating the association of alcohol consumption on incident TB infection with the risk ratio: $\Psi(\mathbb{P})=\Psi(\mathbb{P};1)\div\Psi(\mathbb{P};0)$.
3.5 Accounting for Participant Dependence
Finally, we outline an approach to account for dependence between participants within groups or clusters, such as households, schools, hospitals, or communities. Such dependence could arise due to shared exposures and/or the spread of social behaviors or infectious diseases. In our running example, alcohol use may be influenced by members of a participant’s social circle, and TB is transmitted from person to person. This dependence should be reflected in the corresponding causal model (e.g., [41, 42, 43, 44]). Importantly, by following the Causal Roadmap or a similar framework for causal inference [45, 46], we specify causal models encoding our knowledge about the hierarchical data generating process without imposing parametric modeling assumptions — in contrast to more traditional approaches, such as generalizing estimating equations or mixed effects models (e.g., [47, 44, 22, 23, 48]).
Suppose it is reasonable to assume that participants are dependent within households, but households are effectively independent. (We relax this assumption below.) Then our causal model would be specified at the household-level, and identification would consider the influence of other household members as well as community-level factors. Concretely, this may involve including community indicators in $L_{0}$ and summary measures of household-level covariates in $L_{0}$ and $L_{1}$. In other settings, we may need to expand out the causal model and identification results to accommodate arbitrary dependence across a participant’s social network or across participants within a community. The exact form of the causal model and identification result will depend on the application. Going forward, we use “cluster” to refer to any group considered to be the (conditionally) independent unit [47, 23, 48].
4 Statistical Estimation and Inference
In the previous section, we introduced a series of causal models and identification results of increasing complexity. For the resulting statistical estimands, we could use a singly robust estimation approach, such as G-computation or inverse probability weighting (IPW) [12, 11]. Here, we highlight the use of targeted minimum loss-based estimation (TMLE), which is a doubly robust estimation procedure and asymptotically efficient under certain conditions [14]. In TMLE, initial estimates of the relevant pieces of the observed data distribution are updated to achieve the optimal bias-variance trade-off for the estimand and to solve the efficient influence curve equation. Initial estimates are often computed via Super Learner, an ensemble machine learning algorithm using V-fold cross-validation to select an optimal weighted linear combination of predictions from a library of candidate learners [49]. Thereby, TMLE harnesses machine learning to avoid introducing new modeling assumptions during estimation, while supporting valid statistical inference under reasonable conditions. Notably, for ratio-type estimands corresponding to Counterfactual Strata Effects (Eq. 2), we would implement a separate TMLE for the estimand in the numerator (the joint probability) and the estimand in the denominator (the marginal baseline probability) before combining the results.
TMLE is an asymptotically linear estimator and is normally distributed in the large data limit [14]. When the $N$ participants are i.i.d., the estimator minus the estimand behaves like a sample mean in the first order: $\hat{\Psi}-\Psi=\frac{1}{N}\sum_{i=1}^{N}D_{i}+o_{P}(N^{-1/2})$ where $D_{i}$ is the influence curve for participant $i=\{1,\ldots,N\}$ and $o_{P}(N^{-1/2})$ is a second-order remainder term going to zero in probability [50]. The estimated influence curve is used to calculate standard errors, Wald-type confidence intervals, and p-values. Concretely, a 95% confidence interval is constructed using $\hat{\Psi}\pm z_{0.0975}\frac{\hat{\sigma}}{\sqrt{n}}$ where $z_{0.975}$ is the critical value at the 97.5th-percentile of the standard normal and $\hat{\sigma}$ is the standard deviation of the estimated influence curve. For ratio-type estimands (Eq. 2), once the influence curves for the numerator and denominator have been estimated, the Delta method provides an estimate of the influence curve for our overall estimand (i.e., numerator/denominator). Then to calculate measures of association on the difference, ratio, or odds ratio scale, we apply the Delta method a second time to get inference for these types of functionals.
If clustering is present, estimation and inference must be adjusted. First, the cross-validation scheme used within Super Learner must respect the independent unit. Concretely, participants in a given cluster are all assigned to the same sample-split. Second, for influence curve-based inference, let $m=\{1,\ldots,M\}$ index the clusters and $j=\{1,\ldots,Z_{m}\}$ index for participants in cluster $m$ [51]. Then the total number of participants is $N=\sum_{m}Z_{m}$, and the asymptotic linearity result is re-expressed as $\hat{\Psi}-\Psi=\frac{1}{M}\sum_{m=1}^{M}\left(\sum_{j\in Z_{m}}D_{m,j}\frac{M}{N}\right)$ where $D_{m,j}$ denotes the influence curve for the $j^{th}$ participant in the $m^{th}$ cluster and where we suppressed the second-order remainder term for notational convenience. Altogether, $X_{m}=\frac{M}{N}\sum_{j\in Z_{m}}D_{m,j}$ is the cluster-level influence curve, which has aggregated the individual-level influence curves within cluster $m$ and is weighted by the ratio of the number of clusters to the number of individuals $M/N$. We then proceed with variance estimation using the cluster-level influence curve. This approach is equivalent to using an independent working correlation matrix when obtaining robust (sandwich-based) inference.
5 Application to SEARCH-TB
We now return to our motivating question: what is the effect of alcohol use on incident TB infection among adults in rural Eastern Uganda? With our multinational and interdisciplinary team, we worked through the Causal Roadmap to specify the Statistical Analysis Plan, including the causal model and the adjustment sets needed to account for confounding, missingness, and participant dependence [45, 52, 53, 31]. Our adjustment set included the SEARCH trial arm, community indicators, household HIV status, as well as individual-level age, sex, and mobility measures. For the primary analysis, we used TMLE with Super Learner to combine estimates from generalized linear models, multivariate adaptive regression splines, and the mean. We conducted influence curve-based inference, accounting for clustering by houshold [47, 23]. In secondary analyses, we considered communities, instead of households, to be the independent unit. To examine the impact of modeling assumptions, we also implemented the inverse probability weighting estimator (IPW) with the same adjustment set, but using parametric regressions to estimate the weights. Unlike TMLE, IPW is singly robust — relying on consistent estimation of the propensity score regressions — and tends to be inefficient [14]. Finally, to examine the impact of our missing data assumptions, we implemented the complete case analysis: the simple ratio of mean outcomes after restricting to participants whose exposure and outcomes were measured. For statistical inference, both IPW and the complete case approach accounted for clustering by household.
In the primary analysis (using TMLE with clustering by household), we found that alcohol use was associated with a 49% increase in the risk of incident TB, after flexibly accounting for confounding and missing data: risk ratio (RR)=1.49 (95%CI: 1.39-1.59) [31]. As shown in Figure 5, secondary analyses with the community as the independent unit yielded very similar results, despite meaningfully reducing the effective sample size from 1,380 households to 9 communities: RR=1.49 (95%CI: 1.37-1.62). In contrast, IPW relying on parametric assumptions resulted in a smaller association and very wide confidence intervals: RR=1.13 (95%CI: 1.00-1.27). Finally, after restricting to participants who responded to questions about alcohol use, tested negative at baseline, and tested again at follow-up, the complete case analysis was the least precise and resulted in very wide confidence intervals: RR=1.18 (95%CI: 0.89-1.57).
Figure 5: Results from SEARCH-TB for the association of alcohol use on incident tuberculosis (TB) infection: “Primary” with TMLE and clustering by household, “Secondary” with TMLE and clustering by community, “IPW” with inverse probability weighting, and “Complete Case” with an unadjusted estimator after excluding participants with missing exposures and outcomes.
6 Discussion
We presented causal models, causal parameters, and identification results for a series of observational studies with increasing levels of missingness. For estimation and inference, we highlighted the use of TMLE with Super Learner to robustly and efficiently estimate the corresponding statistical estimands. Application to real-data from SEARCH-TB demonstrated the real-world consequences of our work. Using TMLE to flexibly account for confounding, missingness, and dependence, we found a 49% relative increase in the risk of incident TB associated with drinking alcohol: RR=1.49 (95%CI: 1.39-1.59). With the same adjustment approach but using parametric regressions, IPW resulted in a smaller association and meaningfully wider confidence intervals: RR=1.13 (95%CI: 1.00-1.27). Finally, without adjustment, the complete case analysis yielded a null association (RR=1.18, 95%CI: 0.89-1.57).
There are limitations to our work. First, we did not provide an exhaustive set of causal models and identification results for all possible studies; however, our approach is generalizable and covers many scenarios arising in public health. Second, in our real-data application, we did not include multiple imputation, which is a common approach for missing data and can also leverage machine learning [54, 55, 56]. Future work is needed to investigate the assumptions, implementation, and performance of MI in settings mirroring our motivating example: (1) missingness on the exposure of interest, (2) missingness on the baseline outcome, crucial to defining the target population, (3) missingness in the final outcome, (4) confounding, and (5) dependence among study participants. Third, we relied on various versions of the missing-at-random assumption throughout. In practice, data may be missing as a result of unobserved variables. When data are missing-not-at-random, we may need to collect additional data and conduct sensitivity/bias analyses [57, 10]. Nonetheless, even when a “causal gap” remains, we have a framework to define a statistical estimand, which is aligned with our research question [45, 58, 52, 53]. Finally, we note that interpretation of causal estimands is nuanced when the exposure influences measurement and outcomes (Figures 3 and 4). By considering hypothetical interventions to ensure outcome measurement, we are blocking part of the exposure’s effect. In other words, our causal estimands correspond to a controlled direct effect — not the total effect. In a competing events setting, Young and colleagues have made similar points and defined causal estimands in terms of “separable” direct and indirect effects [15, 59, 60].
Overall, our goal was to contribute to the missing data literature by providing a framework to avoid complete case analyses and, instead, transparently state the missing data assumptions and robustly estimate the corresponding statistical estimands. To do so, we offered guidance for defining and identifying causal effects in real-world studies with missing exposures, missing outcomes, and dependence. We also demonstrated the real-world consequences of our causal and statistical assumptions in SEARCH-TB.
Source of Funding: This work was supported, in part, by The National Institutes of Health (awards: R01AI151209 (CM), K23AI118592 (CM), U01AI099959, and UM1AI068636), the President’s Emergency Plan for AIDS, and the AIDS Research Institute at the University of California San Francisco.
Data and Code Availability: A de-identified dataset and computing code sufficient to reproduce the study findings will be made available following approval of a concept sheet summarizing the analyses to be done. Further inquiries can be directed to the SEARCH Scientific Committee at douglas.black@ucsf.edu.
Acknowledgments: We thank the Ministries of Health of Uganda and Kenya; our research and administrative teams in San Francisco, Uganda, and Kenya; our collaborators and advisory boards; and, especially, all the communities and participants involved. We also thank Dr. Diane Havlir and Dr. Maya L. Petersen, who together with Dr. Moses R. Kamya are the MPIs of the SEARCH collaboration.
7 Supplementary Digital Content for “Causal Inference with Missing Exposures, Missing Outcomes, and Dependence”
In the following, we provide proofs for the identification results. To match the applied example, we focus on binary outcomes, but our results generalize to all outcome-types. For simplicity we focus on categorical covariates, but our summations generalize to integrals for continuous covariates.
eAppendix A: Missing exposures (Figure 2 in the main text)
Let $Y^{*}=Y^{\Delta_{A}=1,A=a}$. Then we have equivalence between our wished-for causal estimand and the corresponding statistical estimand under the following identifability assumptions:
$\mathbb{P}(Y^{*}=1)$ $=$ $\sum_{l}\mathbb{P}(Y^{*}=1\mid L=l)\mathbb{P}(L=l)$
$\text{by }Y^{*}\perp\Delta_{A}\mid L$
$=$ $\sum_{l}\mathbb{P}(Y^{*}=1\mid\Delta_{A}=1,L=l)\mathbb{P}(L=l)$
$\text{by }Y^{*}\perp A\mid\Delta_{A}=1,L$
$=$ $\sum_{l}\mathbb{P}(Y=1\mid A=a,\Delta_{A}=1,L=l)\mathbb{P}(L=l)$
$=$ $\mathbb{E}\big{[}\mathbb{E}(Y\mid A=a,\Delta_{A}=1,L)\big{]}$
For the corresponding statistical estimand to be well-defined, we also need the following positivity assumptions: $\mathbb{P}(\Delta_{A}=1|L)>0$ a.e. and $\mathbb{P}(A\mid\Delta_{A}=1,L)>0$ a.e..
eAppendix B: Missing Exposures and Outcomes (Figure 3 in the main text)
Let $Y^{*}=Y^{\Delta_{A}=1,A=a,\Delta_{Y}=1}$. Then we have equivalence between our wished-for causal estimand and the corresponding statistical estimand under the following identifability assumptions:
$\mathbb{P}(Y^{*}=1)$ $=$ $\sum_{l_{0}}\mathbb{P}(Y^{*}=1\mid L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$\text{by }Y^{*}\perp\Delta_{A}\mid L_{0}\text{ and }Y^{*}\perp A\mid\Delta_{A}=1,L_{0}$
$=$ $\sum_{l_{0}}\mathbb{P}(Y^{*}=1\mid\Delta_{A}=1,A=a,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$\text{by }Y^{*}\perp\Delta_{Y}\mid L_{1},A=a,\Delta_{A}=1,L_{0}$
$=$ $\sum_{l_{0}}\sum_{l_{1}}\mathbb{P}(Y^{*}=1\mid\Delta_{Y}=1,L_{1}=l_{1},\Delta_{A}=1,A=a,L_{0}=l_{0})\times$
$\quad\mathbb{P}(L_{1}=l_{1}\mid\Delta_{a}=1,A=a,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$=$ $\mathbb{E}\bigg{\{}\mathbb{E}\big{[}\mathbb{E}(Y\mid\Delta_{Y}=1,L_{1},A=a,\Delta_{A}=1,L_{0})\mid A=a,\Delta_{A}=1,L_{0})\big{]}\bigg{\}}$
where the inner expectation averages out the outcome $Y$ given the conditioning set, the middle expectation average out the time-varying covariates $L_{1}$ given the conditioning set, and the outer expectation averages out the baseline covariates $L_{0}$.
For the corresponding statistical estimand to be well-defined, we also need the following positivity assumptions: $\mathbb{P}(\Delta_{Y}=1\mid L_{1},A=a,\Delta_{A}=1,L_{0})>0$ a.e. in addition to the positivity assumption from the previous section.
eAppendix C: Missing Exposures and Outcomes at Baseline and Follow-up (Figure 4 in the main text)
Let $Y_{0}^{*}=Y_{0}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1}$ and $Y_{1}^{*}=Y_{1}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1,\Delta_{Y_{1}}=1}$. Recall that we defined the target parameter for this section as
$\mathbb{P}(Y_{1}^{*}=1\mid Y_{0}^{*}=0)=\frac{\mathbb{P}(Y_{1}^{*}=1,Y_{0}^{*}=0)}{\mathbb{P}(Y_{0}^{*}=0)}$
Using the form of the target parameter on the right-hand side of the above equation, we proceed by presenting a separate identification result for the numerator and denominator separately.
C.1 Identification proof for the denominator
Under the following assumptions, which are analogous to eAppendix A, we can identify 1 minus the denominator:
$\mathbb{P}(Y_{0}^{*}=1)$ $=$ $\sum_{l_{0}}\mathbb{P}(Y_{0}^{*}=1\mid L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$\text{by }Y_{0}^{*}\perp\Delta_{A}\mid L_{0}\text{ and }Y_{0}^{*}\perp A\mid\Delta_{A}=1,L_{0}\text{ and }Y_{0}^{*}\perp\Delta_{Y_{0}}\mid A=a,\Delta_{a}=1,L_{0}$
$=$ $\sum_{l_{0}}\mathbb{P}(Y_{0}^{*}=1\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$=$ $\mathbb{E}\big{[}\mathbb{E}(Y\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\big{]}$
along with the corresponding positivity assumptions.
C.2 Identification proof for the numerator
Let $Z^{*}=\mathbb{I}(Y_{1}^{*}=1,Y_{0}^{*}=0)$. Then under the following assumptions, we can identify the numerator $\mathbb{P}(Y_{1}^{*}=1,Y_{0}^{*}=0)=\mathbb{P}(Z^{*}=1)$.
$\mathbb{P}(Z^{*}=1)$ $=$ $\sum_{l_{0}}\mathbb{P}(Z^{*}=1\mid L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$\text{by }Z^{*}\perp\Delta_{A}\mid L_{0}\text{ and }Z^{*}\perp A\mid\Delta_{A}=1,L_{0}\text{ and }Z^{*}\perp\Delta_{Y_{0}}\mid A=a,\Delta_{A}=1,L_{0}$
$=$ $\sum_{l_{0}}\mathbb{P}(Z^{*}=1\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$=$ $\sum_{l_{0}}\sum_{y_{0}}\sum_{l_{1}}\mathbb{P}(Z^{*}=1\mid L_{1}=l_{1},Y_{0}=y_{0},\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\times$
$\quad\mathbb{P}(L_{1}=l_{1},Y_{0}=y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
by $\text{by }Z^{*}=0\text{ when }Y_{0}=1$
$=$ $\sum_{l_{0}}\sum_{l_{1}}\mathbb{P}(Z^{*}=1\mid L_{1}=l_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\times$
$\quad\mathbb{P}(L_{1}=l_{1},Y_{0}=0\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$\text{by }Z^{*}\perp\Delta_{Y_{1}}\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}$
$=$ $\sum_{l_{0}}\sum_{l_{1}}\mathbb{P}(Z^{*}=1\mid\Delta_{Y_{1}}=1,L_{1}=l_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\times$
$\quad\mathbb{P}(L_{1}=l_{1},Y_{0}=0\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$=$ $\sum_{l_{0}}\sum_{l_{1}}\mathbb{P}(Z^{*}=1\mid\Delta_{Y_{1}}=1,L_{1}=l_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\times$
$\quad\mathbb{P}(L_{1}=l_{1}\mid Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\times$
$\quad\mathbb{P}(Y_{0}=0\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$=$ $\mathbb{E}[\mathbb{E}\left(\mathbb{E}(Y_{1}\mid\Delta_{Y_{1}}=1,L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}\right)]$
For the corresponding statistical estimand to be well-defined, we also need the following positivity assumptions: $P(\Delta_{Y_{1}}=1\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})>0$ a.e. in addition to the positivity assumption for the denominator.