### v2

# Causal Inference with Missing Exposures and Missing Outcomes

Abstract
Missing data are ubiquitous in public health research. When estimating causal effects, there are well-established methods to address bias to due censored outcomes. Commonly, causal estimands are defined under hypothetical interventions to “set” the exposure and to prevent censoring. Identification is evaluated with the sequential backdoor criterion and considerations of data support. Then inverse weighting, standardization, and doubly-robust approaches are applied for statistical estimation and inference. We demonstrate how this framework can be extended to settings with missingness on the exposure of interest as well as the variable defining the population of interest (e.g., persons at risk of the outcome). Our work is motivated by SEARCH-TB’s investigation of the effect of alcohol consumption on the risk of incident tuberculosis (TB) infection in rural Uganda. This study posed several real-world challenges: confounding, missingness on the exposure (alcohol use), missingness on the baseline outcome (defining who was at risk of TB), and missingness on the outcome at follow-up (capturing who acquired TB). We present a series of causal models and identification results to demonstrate the handling of missing exposures and outcomes in prospective studies. We highlight the use of TMLE with Super Learner and the real-world consequences of our approach.
Keywords: Causal Inference, Missing Data, Real-world Evidence, Super Learner, Targeted Minimum Loss-based Estimation, TMLE, Tuberculosis
Introduction
Missing data affect the integrity of analyses across the spectrum of public health research, including surveillance studies to estimate disease prevalence and randomized trials to establish efficacy of new medical products [1, 2, 3, 4, 5, 6, 7, 8]. There is rich history of methods research to address the potential for bias when participants with measured outcomes differ meaningfully from those with missing or censored outcomes (e.g., [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]). The Causal Roadmap provides one such approach [21, 22, 23]. First, we specify our causal question: how would expected outcomes differ if all did versus did not receive the exposure of interest and censoring were prevented. Then, we specify a causal model, such as directed acyclic graph (DAG) or non-parametric structural equation model (NPSEM), to represent the data generating process for observed data: baseline and time-varying confounders, exposures, and outcomes [10]. Third, we intervene on the causal model to generate counterfactual outcomes under hypothetical interventions to set the exposures and ensure outcome measurement. Fourth, we evaluate whether the corresponding causal estimand can be identified (i.e., if sequential exchangeability and positivity hold) [9, 10, 11, 12]. Fifth, we specify the statistical estimand, which is often a complex function of the observed data distribution (i.e., not equal to single regression coefficient). Sixth, we conduct estimation and inference with inverse weighting, standardization, or doubly-robust methods, such as targeted minimum loss-based estimation (TMLE) [11]. Finally, we conduct sensitivity analyses and appropriately interpret the results.
Here, we provide a tutorial on how this framework can be extended to address missingness on the exposure. We also highlight the consequences of missingness on the baseline outcome when it is crucial to defining the population of interest. Suppose, for example, we are interested in studying the incidence of some disease. Our population of interest would be persons who are at risk of the developing the outcome and are, thereby, disease-free at baseline. In this setting, our incidence estimates would be subject to bias if there is differential missingness of outcomes at baseline. Using Counterfactual Strata Effects [24, 5, 25, 26, 27, 28, 29], we provide a framework for explicitly defining, identifying, and estimating parameters in such scenarios.
We note there is a growing interest in the use of missingness graphs to represent studies with missingness on multiple variables and to assess whether causal effects can be identified (termed “recovered”) [30, 31]. In particular, Moreno-Betancur and colleagues introduced complete-data DAGs (c-DAGs), missingness DAGs (m-DAGs), and their link [4]. They also provided a series of “canonical” m-DAGs for causal effects of point-treatment exposures [4, 32]. Holovchak et al. extended this work for the effects of longitudinal exposures [33], while questioning the practical utility of the m-DAG approach. Specifically, they noted, “no general algorithms are available to decide on recoverability, and decisions have to be made on a case-by-case basis” [33]. As an alternative, we demonstrate how an established framework for causal inference can used to define, identify, and estimate causal estimands, such as the average treatment effect, under missingness on exposures and outcomes. In the following, we build-up from simple to complex examples in hope that our structured presentation is relevant a wide range of readers who need to address confounding and multi-source missingness. We illustrate practical relevance with the SEARCH-TB study.
Motivating example
SEARCH was a cluster randomized trial to evaluate a community-based approach to a Universal HIV Test-and-Treat intervention, as compared to the standard-of-care, in rural Kenya and Uganda (2013-2017; NCT01864603) [34]. Following a rapid census, all communities were offered multi-disease testing through community health campaigns with home-based follow-up for non-participants [35]. Through this mechanism, we measured demographic data (e.g., age, sex, education, and mobility), self-reported alcohol use (our primary exposure of interest), and tested for HIV infection. Due to high costs and complex logistics, evaluation of incident tuberculosis (TB) infection was limited to SEARCH-TB, a sub-study in 9 eastern Ugandan communities [36, 37]. This sub-study was intentionally enriched for persons with HIV. Specifically, in each community, we sampled 100 households with at least one adult (15+ years) with HIV and 100 households without an adult with HIV. At baseline of the sub-study, tuberculin skin tests were administered to residents of the sampled households. One year later, follow-up tests were administered to participants who tested negative at baseline. Here, we demonstrate the methods used to evaluate the effect of alcohol use on incident TB infection [38]: confounding, missingness on the exposure of interest, missingness on the baseline outcome (defining who was at risk of TB), and missingness on the final outcome (defining who acquired TB).
Related Causal Problems: Building Complexity
Many studies feature only a subset of the challenges described above. We, thus, provide causal models and identification results for a series of hypothetical studies with increasing complexity in the hope of providing a useful reference for a broad range of real-world studies. For simplicity, we focus on defining and identifying causal parameters under a single level of the exposure, but our results naturally generalize to causal effects defined in terms of contrasts of counterfactual outcome distributions under two levels of the exposure (i.e., the average treatment effect or causal risk ratio).
Classic point-treatment problem
In Appendix S1, we review the classic “point-treatment” problem, where we have measured confounding by baseline covariates $L$, a binary exposure $A$ occurring at single time-point, and an outcome $Y$ occurring at the study’s close. This could represent a study of the effect of alcohol use ($A$) on incident TB infection ($Y$) among a representative cohort of persons without TB at baseline. Let $Y^{a}$ be the counterfactual outcome if, possibly contrary-to-fact, the participant had exposure-level $A=a$. The counterfactual mean outcome $\mathbb{E}(Y^{a})$ is identified if there are no unmeasured confounders and sufficient data support: $Y^{a}\perp A\mid L$ and $\mathbb{P}(A=a\mid L)>0$, respectively. The statistical estimand is given by the G-computation formula: $\mathbb{E}[\mathbb{E}(Y\mid A=a,L)]$[9]. Even if these assumptions are not reasonable, the G-computation formula is still a well-defined statistical estimand, on which we can focus our estimation efforts (Appendix S1).
Missing Outcomes
Most real-world studies, however, depart from this idealized point-treatment problem. Here, we review causal models and identification when there is missingness on the outcome. Continuing our running example, suppose that among our representative cohort of persons without TB at baseline, some participants did not test for TB at the end of follow-up. Let $\Delta_{Y}$ be an indicator of outcome measurement. If $\Delta_{Y}=1$ for a participant, we observe their outcome $Y$ as usual. However, if $\Delta_{Y}=0$, their outcome $Y$ is not observed. In Figure 1, we show two possible ways of using causal models to represent such a study. In Panel A, we introduce $Y^{\circ}$ to represent the underlying value of the outcome, and define the observed outcome as the product of its measurement indicator and underlying value: $Y=\Delta_{Y}\times Y^{\circ}$. Thus, $Y$ equals its underlying value $Y^{\circ}$ when it’s measured (i.e., when $\Delta_{Y}=1$) and is zero otherwise (i.e., when $\Delta_{Y}=0$). In Panel B, we omit the underlying outcome and directly represent the causal model in terms of the observed data: $O=(L,A,\Delta_{Y},Y)$. In the latter presentation, the measurement indicator $\Delta_{Y}$ again influences the value of the observed outcome $Y$, because when an outcome is not measured, it is set deterministically to NA. Both approaches lead us to same identification assumptions and statistical estimands. Therefore, to minimize notation and mirror the long-standing literature on censoring (e.g., [16, 17, 11, 18, 19, 20, 12]), we use the latter representation in the remainder of the article.
To define causal effects when the outcome is subject to missingness, we now consider counterfactuals indexed by both the exposure and the outcome measurement indicator. Specifically, let $Y^{*}=Y^{A=a,\Delta_{Y}=1}$ be the counterfactual outcome for a given participant if, possibly contrary-to-fact, their exposure were at level $A=a$ and missingness on the outcome prevented. Then our causal target parameter $\mathbb{E}(Y^{*})$ is the counterfactual mean outcome if all participants had exposure $A=a$ and their outcomes measured. To identify this causal parameter and express it as a function of the observed data distribution, we would need the baseline covariates $L$ to be sufficient to control for confounding and differential outcome missingness. Specifically, we need $L$ to capture all the common causes of the exposure and outcome, and, among those with the exposure of interest, all the common causes of the outcome and its measurement. These assumptions can be represented as $Y^{*}\perp A\mid L$ and $Y^{*}\perp\Delta_{Y}\mid A=a,L$, respectively. These conditions are often referred to as the “sequential randomization assumption” or “sequential exchangeability” and can be evaluated graphically through the sequential backdoor criterion [9, 10, 11, 12]. We also need a positive probability of being exposed to level $A=a$ within all possible values of the confounders and, among those with the exposure of interest, a positive probability of outcome measurement within all possible values of the confounders: $\mathbb{P}(A=a\mid L)>0$ a.e. and $\mathbb{P}(\Delta_{Y}\mid A=a,L)>0$ a.e., respectively. If these assumptions hold, we have equivalence between our causal estimand $\mathbb{E}(Y^{*})$ and the statistical estimand given by the G-computation formula: $\mathbb{E}[\mathbb{E}(Y\mid\Delta_{Y}=1,A=a,L)]$ [9].
Figure 1: Causal models with missingness on the outcome (Panels A-B) or missingness on the exposure (Panel C): LL=baseline covariates, $\Delta_{A}$=indicator of exposure measurement, AA=observed exposure, $Y^{\circ}$=underlying outcome, $\Delta_{Y}$=indicator of outcome measurement, YY= observed outcome. For ease of presentation, the models are shown without dependence between the unmeasured variables, which are omitted on the graph.
Missing Exposures
We now demonstrate how the above approach for missing outcomes can be applied to missing exposures. Continuing our running example, suppose that among our representative cohort of persons without TB at baseline, some participants did not answer questions about their alcohol use. Let $\Delta_{A}$ be an indicator that a participant has their exposure measured. If $\Delta_{A}=1$ for a participant, we observe their exposure $A$ as usual. However, if $\Delta_{A}=0$ for a participant, their exposure $A$ is set deterministically to NA. The DAG and NPSEM for such a study are given in Figure 1C. To define causal effects when the exposure is subject to missingness, we consider counterfactuals indexed by both the exposure and its measurement indicator. Now, let $Y^{*}=Y^{\Delta_{A}=1,A=a}$ be the counterfactual outcome for a given participant if, possibly contrary-to-fact, their exposure were measured and at level $A=a$.
Then to identify the expected counterfactual outcome $\mathbb{E}(Y^{*})$ and express it as a function of the distribution of the observed data $O=(L,\Delta_{A},A,Y)$, we need analogous conditions as the prior subsection on missing outcomes. Specifically, we need $L$ to be sufficient to control for differential exposure missingness and for confounding (among those with measured exposures): $Y^{*}\perp\Delta_{A}\mid L$ and $Y^{*}\perp A\mid\Delta_{A}=1,L$, respectively. These assumptions can be graphically evaluated with the sequential backdoor criterion. We also need the two analogous assumptions on data support: $\mathbb{P}(\Delta_{A}=1\mid L)>0$ a.e. and $\mathbb{P}(A=a\mid\Delta_{A}=1,L)>0$ a.e. If these four assumptions hold, we can rewrite $\mathbb{E}(Y^{*})$ as the statistical estimand $\mathbb{E}[\mathbb{E}(Y\mid A=a,\Delta_{A}=1,L)]$ with proof in Appendix S2.1.
Missing Exposures and Outcomes
We now combine our two challenges: missing exposures and missing outcomes. Continuing our running example, suppose that among our cohort of persons at risk of TB, some participants did not answer questions about their alcohol use and, despite best efforts, some participants could not be found at the end of the study for outcome ascertainment. To reflect this data generating process, we introduce new notation to reflect the longitudinal data setting. Let $L_{0}$ be baseline covariates and $L_{1}$ be additional covariates collected after the exposure but before outcome ascertainment. Figure 2 provides the causal models for such a study.
Figure 2: Causal graph and corresponding non-parametric structural equation model with missingness on the exposure and the outcome: $L_{0}$=baseline covariates, $\Delta_{A}$=indicator of exposure measurement, AA=observed exposure, $L_{1}$=time-varying covariates, $\Delta_{Y}$=indicator of outcome measurement, and YY=observed outcome. For ease of presentation, the models are shown without dependence between the unmeasured variables, which are omitted on the graph.
To define the causal effect when the exposure and outcome are subject to missingness, we now consider counterfactuals indexed by the exposure and two measurement indicators. Specifically, let $Y^{*}=Y^{\Delta_{A}=1,A=a,\Delta_{Y}=1}$ denote the counterfactual outcome under hypothetical interventions to ensure exposure measurement, “set” the exposure level to $A=a$, and ensure outcome measurement. To identify the counterfactual mean outcome $\mathbb{E}(Y^{*})$ and express it as a function of the distribution of the observed data $O=(L_{0},\Delta_{A},A,L_{1},\Delta_{Y},Y)$, we now need to account for the post-baseline covariates $L_{1}$, which act as time-dependent confounders. Specifically, $L_{1}$ are mediators of the exposure-outcome relationship, while “confounding” the measurement-outcome relationship. Therefore, we rely on sequential randomization/exchangeability and find a set of covariates that satisfies the backdoor criterion for each “intervention” node given the observed past [9]. As before, we need that the baseline covariates $L_{0}$ are sufficient to control for missing exposures and for confounding. In other words, we need the analogous identification assumptions given in the prior subsection. Additionally, we need that among participants with measured exposures at the level of interest (i.e., $\Delta_{A}=1$ and $A=a$), the baseline and time-varying covariates ($L_{0},L_{1})$ capture all the common causes of outcomes and their measurement as well as a positive probability of outcome measurement within all possible values of the baseline and time-varying covariates: $Y^{*}\perp\Delta_{Y}\mid L_{1},A=a,\Delta_{A}=1,L_{0}$ and $\mathbb{P}(\Delta_{Y}=1\mid L_{1},A=a,\Delta_{A}=1,L_{0})>0$ a.e., respectively. If these assumptions hold, we can rewrite $\mathbb{E}(Y^{*})$ in terms of the longitudinal G-computation formula: $\mathbb{E}\{\mathbb{E}[\mathbb{E}(Y\mid\Delta_{Y}=1,L_{1},A=a,\Delta_{A}=1,L_{0})\mid A=a,\Delta_{A}=1,L_{0})]\}$, shown in terms of iterated expectations and with proof in Appendix S2.2 [9, 39, 40].
Missing Exposures and Missing Outcomes at the Start and End of Follow-up
We now consider missingness on the outcome at the start of follow-up. Following our motivating example, we are interested in the effect of alcohol use on incident TB infection, but did not reach 100% of study participants for baseline TB testing. In other words, our longitudinal cohort of participants without TB is subject to selection bias. To reflect this data generating process, we update our notation to have multiple outcome measures. Let $\Delta_{Y_{0}}$ be an indicator of outcome measurement at start of follow-up (hereafter “baseline”) and $\Delta_{Y_{1}}$ be an indicator of outcome measurement at the end of follow-up (hereafter “endline”). Let $Y_{0}$ and $Y_{1}$ denote the corresponding outcomes; they are set to NA when not observed. The corresponding causal models can be found in Figure 3. The time-ordering reflects the study protocol and procedures in SEARCH-TB; specifically, alcohol use $A$ was measured before baseline TB status $Y_{0}$. As a result, the exposure $A$ can impact who has prevalent TB at baseline ($Y_{0}=1$) and, thereby, who is at risk of TB at baseline ($Y_{0}=0$).
Figure 3: Causal graph and corresponding non-parametric structural equation model with missingness on the exposure and the outcome at baseline and endline: $L_{0}$=baseline covariates, $\Delta_{A}$=indicator of exposure measurement, A=exposure, $\Delta_{Y_{0}}$=indicator of outcome measurement at baseline, $Y_{0}$=baseline outcome, $L_{1}$=time-dependent covariates, $\Delta_{Y_{1}}$=indicator of outcome measurement at endline, and $Y_{1}$= outcome at endline. For ease of presentation, the models are shown without dependence between the unmeasured variables, which are omitted on the graph.
To define the causal target parameter in this setting, we first consider the counterfactual outcome at baseline under hypothetical interventions to ensure exposure measurement, “set” the exposure level to $A=a$, and ensure outcome measurement at baseline: $Y_{0}^{*}=Y_{0}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1}$. Additionally, we consider the counterfactual outcome at endline under the prior interventions as well as a hypothetical intervention to ensure outcome measurement at endline among those known to be at risk at baseline. Concretely, only participants who tested TB-negative at baseline ($\Delta_{Y_{0}}=1$ and $Y_{0}=0$) were approached for re-testing at endline. Thereby, this final intervention is a dynamic or personalized one to set $\Delta_{Y_{1}}$ equal to one if $Y_{0}=0$ and to zero otherwise (e.g., [24, 41, 42, 43]). For simplicity, denote the resulting counterfactual outcome as $Y_{1}^{*}=Y_{1}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1,\Delta_{Y_{1}}=1}$.
Now we can precisely define the causal parameter in terms of the following conditional probability, which captures the counterfactual incidence of the outcome among those at risk at baseline: $\mathbb{P}(Y_{1}^{*}=1\mid Y_{0}^{*}=0)$. Due to conditioning on a counterfactual variable, such parameters are sometimes called “Counterfactual Strata Effects”[24, 5, 25, 26, 27, 28], which are defined by Nakato et al. as “causal estimands where the outcome is only relevant for a group whose membership is subject to missingness and/or impacted by the exposure” [29]. These effects are different from Principal Strata Effects, which are defined within subgroups of latent classes that are fundamentally not observable [44, 45, 46, 47]. For example, in SEARCH-TB, principal stratification could be applied to define the effect of alcohol use on incident TB infection among the subset of participants who would have always tested regardless of their alcohol use. Instead, our interest is the effect of alcohol use on incident TB among the entire population of persons at risk at baseline.
To identify this effect, we re-express the conditional probability as
$\mathbb{P}(Y_{1}^{*}=1\mid Y_{0}^{*}=0)=\frac{\mathbb{P}(Y_{1}^{*}=1,Y_{0}^{*}=0)}{\mathbb{P}(Y_{0}^{*}=0)}$ (1)
Then given the observed data $O=(L_{0},\Delta_{A},A,\Delta_{Y_{0}},Y_{0},L_{1},\Delta_{Y_{1}},Y_{1})$, we can identify denominator and numerator, in turn. The denominator $\mathbb{P}(Y_{0}^{*}=0)$ represents the counterfactual prevalence of not having the outcome at baseline and, thus, being at risk. The causal structure for this parameter is analogous to that of Section 3.2, but with an additional measurement indicator for the baseline outcome. Therefore, under analogous assumptions, we can identify $\mathbb{E}(Y_{0}^{*})=\mathbb{E}\big[\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\big]$ with proof in Appendix S2.3. Since we are interested the counterfactual probability of being at risk at baseline $\mathbb{P}(Y_{0}^{*}=0)$, our statistical estimand for the denominator becomes $1-\mathbb{E}\big[\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\big]$.
In our final causal parameter (Eq. 1), the numerator $\mathbb{P}(Y_{1}^{*}=1,Y_{0}^{*}=0)$ represents the counterfactual probability of having the outcome at endline but not at baseline. For ease of notation, let $Z^{*}=\mathbb{I}(Y_{1}^{*}=1,Y_{0}^{*}=0)$ represent the joint indicator of these two counterfactual values. To identify $\mathbb{E}(Z^{*})=\mathbb{P}(Z^{*}=1)$, we need analogous assumptions as for the denominator together with the following. Among those known to be at risk at baseline ($\Delta_{Y_{0}}=1,Y_{0}=0$) and with measured exposure of interest ($\Delta_{A}=1,A=a$): the baseline and time-varying covariates capture the common causes of the joint outcome and endline measurement: $Z^{*}\perp\Delta_{Y_{1}}\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}$. We also need there to be a positive probability of follow-up measurement within all possible values of $L_{0}$ and $L_{1}$: $\mathbb{P}(\Delta_{Y_{1}}=1\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})>0$ a.e. Under these assumptions and with proof given in Appendix S2.3, the numerator is identified as
$\mathbb{P}(Z^{*}=1)\ =\ \mathbb{E}[\mathbb{E}\left(\mathbb{E}(Y_{1}\mid\Delta_{Y_{1}}=1,L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}\right)]$
Putting it all together, the statistical estimand with missing exposures, missing outcomes at baseline, and missing outcomes at endline is given by
$\Psi(\mathbb{P};a)=\frac{\mathbb{E}[\mathbb{E}\left(\mathbb{E}(Y_{1}\mid\Delta_{Y_{1}}=1,L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}\right)]}{1-\mathbb{E}\left[\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\right]}$ (2)
for exposure level $A=a$. Then we can define associations in terms of contrasts $\Psi(\mathbb{P};a)$ at different exposure levels. Concretely, in SEARCH-TB, we were interested in evaluating the association of alcohol consumption on incident TB infection with the risk ratio: $\Psi(\mathbb{P})=\Psi(\mathbb{P};1)\div\Psi(\mathbb{P};0)$.
Statistical Estimation and Inference
In the previous section, we introduced a series of causal models and identification results of increasing complexity. For the resulting statistical estimands, we could use a singly robust estimation approach, such as standardization (a.k.a., “G-computation”) or inverse probability weighting (IPW) [9, 13]. Here, we highlight the use of TMLE, which is a doubly robust estimation procedure and asymptotically efficient under certain conditions [11]. In TMLE, initial estimates of the relevant pieces of the observed data distribution are updated to achieve the optimal bias-variance trade-off for the estimand and to solve the efficient influence curve equation. Initial estimates are often computed via Super Learner, an ensemble machine learning algorithm using V-fold cross-validation to select an optimal weighted linear combination of predictions from a library of candidate learners [48]. Thereby, TMLE leverages machine learning to avoid introducing new modeling assumptions during estimation, while supporting valid statistical inference under reasonable conditions. Notably, for ratio-type estimands corresponding to Counterfactual Strata Effects (Eq. 2), we would implement a separate TMLE for the estimand in the numerator (the joint probability) and the estimand in the denominator (the marginal baseline probability) before combining the results.
TMLE is an asymptotically linear estimator and is normally distributed in the large data limit [11]. The estimator minus the estimand behaves like a sample mean in the first order: $\hat{\Psi}-\Psi=\frac{1}{N}\sum_{i=1}^{N}D_{i}+o_{P}(N^{-1/2})$ where $D_{i}$ is the influence curve for participant $i=\{1,\ldots,N\}$ and $o_{P}(N^{-1/2})$ is a second-order remainder term going to zero in probability [49]. The estimated influence curve is used to calculate standard errors, Wald-type confidence intervals, and p-values. Concretely, a 95% confidence interval is constructed using $\hat{\Psi}\pm z_{0.0975}\frac{\hat{\sigma}}{\sqrt{n}}$ where $z_{0.975}$ is the critical value at the 97.5th-percentile of the standard normal and $\hat{\sigma}$ is the standard deviation of the estimated influence curve. For ratio-type estimands (Eq. 2), once the influence curves for the numerator and denominator have been estimated, the Delta method provides an estimate of the influence curve for our overall estimand. Then to calculate measures of association on the difference, ratio, or odds ratio scale, we apply the Delta method a second time to get inference for these types of functionals.
Application to SEARCH-TB
We now return to our motivating question: what is the effect of alcohol use on incident TB infection among adults in rural Eastern Uganda? With our multinational and interdisciplinary team, we worked through the Causal Roadmap to specify the Statistical Analysis Plan, including the causal model and the adjustment sets [21, 22, 23, 50, 38]. Concretely, the causal model reflected team’s knowledge of study protocol, the study procedures, and the epidemiology of TB in the region. Our adjustment set included the SEARCH trial arm, community indicators, household HIV status, as well as individual-level age, sex, and mobility measures. For the primary analysis, we used TMLE with Super Learner to combine estimates from generalized linear models, multivariate adaptive regression splines, and the mean. We conducted influence curve-based inference, accounting for clustering by household (Appendix S4) [51, 26]. In secondary analyses, we considered communities, instead of households, to be the independent unit. To examine the sensitivity of our results to alternative estimation approaches, we also implemented IPW for the same statistical estimand, but using parametric regressions to estimate the weights. Finally, to examine the impact of handling of our missing data, we took the following “naïve” approach: (1) subset on participants known to be at risk of TB at baseline ($Y_{0}=0$); (2) further subset on participants with measured exposures ($\Delta_{A}=1$) and measured outcomes at endline ($\Delta_{Y_{1}}=1$), and (3) implement TMLE to adjust for confounding and to estimate $\mathbb{E}[\mathbb{E}(Y_{1}\mid A=a,L)\mid Y_{0}=0]$. This approach is inherently flawed, because of the bias induced by inappropriately conditioning on various colliders and mediators (Figure 3). Nonetheless, the approach could be one taken by an analyst aiming to implement a “complete-case” analysis while adjusting for confounding [52, 53, 54]. For statistical inference, both IPW and the naïve approach accounted for clustering by household.
In the primary analysis using TMLE with clustering by household, we found that alcohol use was associated with a 49% increase in the risk of incident TB: risk ratio (RR)=1.49 (95%CI: 1.39-1.59) [38]. As shown in Figure 4, secondary analyses with the community as the independent unit yielded very similar results, despite meaningfully reducing the effective sample size from 1,380 households to 9 communities: RR=1.49 (95%CI: 1.37-1.62); Appendix S4. In contrast, IPW relying on parametric assumptions resulted in a smaller association and confidence intervals overlapping the null: RR=1.13 (95%CI: 1.00-1.27). Finally, after restricting to participants who tested negative at baseline, responded to questions about alcohol use, and tested again at follow-up, the naïve approach was the least precise and resulted in the widest confidence intervals: RR=1.18 (95%CI: 0.89-1.57).
Figure 4: Results from SEARCH-TB for the association of alcohol use on incident tuberculosis (TB) infection: “Primary” with TMLE and clustering by household, “Secondary” with TMLE and clustering by community, “IPW” with inverse probability weighting, and “Naïve” based on subsetting on those at risk at baseline and with measured exposures and outcomes at endline.
Discussion
We presented causal models, causal parameters, and identification results for a series of prospective studies with increasing levels of missingness. For estimation and inference, we highlighted the use of TMLE with Super Learner to robustly and efficiently estimate the corresponding statistical estimands. Application to real-data from SEARCH-TB demonstrated the real-world consequences of our work. Using TMLE to flexibly account for confounding and missingness, we found a 49% relative increase in the risk of incident TB infection associated with drinking alcohol: RR=1.49 (95%CI: 1.39-1.59). For the same statistical estimand but using parametric regressions, IPW resulted in a smaller association and meaningfully wider confidence intervals: RR=1.13 (95%CI: 1.00-1.27). Finally, the naïve approach to condition on the participants with baseline TB risk and measured exposures/outcomes yielded a null association (RR=1.18, 95%CI: 0.89-1.57).
There are several advantages to our framework for handling missing exposures and outcomes when evaluating the causal effects. Unlike the m-DAG approach which aids in identifying the entire joint distribution of the observed data [4], our approach focuses our efforts on the portions of the observed data distribution relevant for the statistical estimand. Furthermore, our approach uses data on all participants, improving efficiency relative to approaches excluding participants with missing data on the relevant variables (sometimes called an “available-case analysis”) [4]. Finally, our approach leads us to statistical estimands that can be robustly and rigorously estimated with modern methods, such as TMLE.
A major limitation to this work is that we did not address missingness on the confounders. Some approaches include mean/median imputation or including missingness indicators in the adjustment set. For further discussion on missing confounders, we refer to the comprehensive work of Williamson et al. [55]. We also focused on cross-sectional or prospective studies. Thus, we did not cover scenarios where the outcome impacts the measurement of other variables. Such scenarios would arise in case-control studies and have been addressed in prior literature (e.g., [56, 57, 58, 59, 54]).
There are other limitations to our work. First, we did not provide an exhaustive set of causal models and identification results for all possible studies; however, our approach is generalizable and covers many scenarios arising in public health. Second, in our real-data demonstration, we did not include multiple imputation, which is a common approach for missing data and can also leverage machine learning [60, 61, 62]. Future work is needed to investigate the assumptions, implementation, and performance of multiple imputation in settings mirroring our motivating example: (1) missingness on the exposure of interest, (2) missingness on the baseline outcome, (3) missingness in the final outcome, (4) confounding, and (5) dependence among study participants. Finally, we relied on various versions of the sequential exchangeability assumption for both confounding and missingness. In practice, data may be missing as a result of unobserved variables, and we may need to collect additional data as well as conduct sensitivity/bias analyses [63, 64]. Nonetheless, even when a “causal gap” remains, we have a framework to define a statistical estimand, which is aligned with our research question [21, 65, 23, 50].
Overall, our goal was to demonstrate how an established framework for estimating causal effects with censored outcomes could be extended to settings with missingness on the exposure, the outcome at baseline, and the outcome at endline. To do so, we provided a series of causal models with increasing complexity and discussed the identification assumptions for each. For estimation and inference, we focused on TMLE, a doubly-robust approach using machine learning and data from all participants (including those with missing exposures or outcomes). Our work was motivated by the investigation of the effect of alcohol use on incident TB infection [38], and our re-analysis demonstrated the real-world consequences of our approach.
Source of Funding: This work was supported, in part, by The National Institutes of Health (awards: R01AI151209 (CM), K23AI118592 (CM), U01AI099959, and UM1AI068636), the President’s Emergency Plan for AIDS, and the AIDS Research Institute at the University of California San Francisco.
Conflicts of Interest: The authors report no conflict of interest in this work.
Acknowledgments: We thank the Ministries of Health of Uganda and Kenya; our research and administrative teams in San Francisco, Uganda, and Kenya; our collaborators and advisory boards; and, especially, all the communities and participants involved. We also thank Dr. Diane Havlir and Dr. Maya L. Petersen, who together with Dr. Moses R. Kamya are the MPIs of the SEARCH collaboration.
Data and Code Availability: A de-identified dataset and computing code sufficient to reproduce the study findings will be made available following approval of a concept sheet summarizing the analyses to be done. Further inquiries can be directed to the SEARCH Scientific Committee at douglas.black@ucsf.edu.
Supplementary Materials for “Causal Inference with Missing Exposures and Missing Outcomes”
Contents:
  • Appendix S1: More on the classic point-treatment problem
  • Appendix S2: Proofs
  • Appendix S3: Accounting for Outcome Dependence
Appendix S1: More on the classic point-treatment problem
We consider the classic “point-treatment” problem, where we have measured confounding by baseline covariates $L$, a binary exposure $A$ occurring at single time-point, and an outcome $Y$ occurring at the study’s close. This could represent a study of the effect of alcohol use ($A$) on incident TB infection ($Y$) among a representative cohort of persons without TB at baseline. The directed acyclic graph (DAG) and non-parametric structural equation model (NPSEM) for such a study are given in Figure 5A.
Under interventions on the causal model, we generate counterfactual outcomes corresponding to the research question of interest. Specifically, let $Y^{a}$ be the counterfactual outcome for a given participant if, possibly contrary-to-fact, they had exposure-level $A=a$. Then our causal target parameter $\mathbb{E}(Y^{a})$ is the counterfactual mean outcome if all study participants had exposure-level $A=a$. In our running example, $\mathbb{E}(Y^{a})=\mathbb{P}(Y^{a}=1)$ is the counterfactual risk of incident TB infection with alcohol use $A=a$.
Figure 5: Causal models for a classic point-treatment problem with complete measurement of the baseline covariates $L$, the exposure $A$, and the outcome $Y$. On the directed acyclic graph, $U^{*}$ represents unmeasured common causes of at least two variables in $(L,A,Y)$. In panel A provides the causal models under no assumptions about the unmeasured factors. Panels B and C are compatible with the no unmeasured confounders assumption.
To identify our causal target parameter and express it as function of the distribution of the observed data $O=(L,A,Y)$, we would need there to be no unmeasured confounding, which corresponds to the assumption that the baseline covariates $L$ capture all the joint causes of the exposure $A$ and outcome $Y$. This condition is called “the randomization assumption” or “exchangeability”, can be evaluated with the backdoor criterion, and can be represented as $Y^{a}\perp A\mid L$ [9, 10, 11, 12].
Concretely, $\mathbb{E}(Y^{a})$ is not identified in Figure 5A because there is an unmeasured common cause, represented by $U^{*}$, of the exposure $A$ and outcome $Y$. In Figure 5B and C, we show two causal models where this assumption would hold.
Additionally, we need there to be a non-zero probability of having the exposure in all possible values of $L$: $\mathbb{P}(A=a\mid L)>0$ a.e. This is a condition on data support and known as the “positivity assumption”. We note that the consistency assumption, stating that the counterfactual outcome $Y^{a}$ equals the observed outcome $Y$ under exposure $A=a$, holds by design when we define counterfactuals through intervention on the causal model. Under the exchangeability and positivity assumptions, our causal target is equal to the statistical estimand given by the G-computation formula: $\mathbb{E}[\mathbb{E}(Y\mid A=a,L)]$ [9]. Even if these assumptions are not reasonable (e.g., Figure 5A reflects reality), we still have well-defined statistical estimand, on which we can focus our estimation efforts. In other words, we still proceed with estimation and inference $\mathbb{E}[\mathbb{E}(Y\mid A=a,L)]$ , while appropriately accounting for lack of identification in our interpretation [21, 22, 23, 66, 65] For ease of presentation, we provide the causal models without dependence between the unmeasured factors in the main text.
Appendix S2: Proofs
In the following, we provide proofs for the identification results. To match the applied example, we focus on binary outcomes, but our results generalize to all outcome-types. For simplicity we focus on categorical covariates, but our summations generalize to integrals for continuous covariates.
Appendix S2.1: Missing exposures (Figure 1C in the main text)
Let $Y^{*}=Y^{\Delta_{A}=1,A=a}$. Then we have equivalence between our wished-for causal estimand and the corresponding statistical estimand under the following identifability assumptions:
$\mathbb{P}(Y^{*}=1)$ $=$ $\sum_{l}\mathbb{P}(Y^{*}=1\mid L=l)\mathbb{P}(L=l)$
$\text{by }Y^{*}\perp\Delta_{A}\mid L$
$=$ $\sum_{l}\mathbb{P}(Y^{*}=1\mid\Delta_{A}=1,L=l)\mathbb{P}(L=l)$
$\text{by }Y^{*}\perp A\mid\Delta_{A}=1,L$
$=$ $\sum_{l}\mathbb{P}(Y^{*}=1\mid A=a,\Delta_{A}=1,L=l)\mathbb{P}(L=l)$
by the consistency assumption
$=$ $\sum_{l}\mathbb{P}(Y=1\mid A=a,\Delta_{A}=1,L=l)\mathbb{P}(L=l)$
$=$ $\mathbb{E}\big[\mathbb{E}(Y\mid A=a,\Delta_{A}=1,L)\big]$
We again note the consistency assumption holds by our definition of counterfactual outcomes as being derived through interventions on the causal model. For the corresponding statistical estimand to be well-defined, we also need the following positivity assumptions: $\mathbb{P}(\Delta_{A}=1|L)>0$ a.e. and $\mathbb{P}(A\mid\Delta_{A}=1,L)>0$ a.e..
Appendix S2.2: Missing Exposures and Outcomes (Figure 2 in the main text)
Let $Y^{*}=Y^{\Delta_{A}=1,A=a,\Delta_{Y}=1}$. Then we have equivalence between our wished-for causal estimand and the corresponding statistical estimand under the following identifability assumptions:
$\mathbb{P}(Y^{*}=1)$ $=$ $\sum_{l_{0}}\mathbb{P}(Y^{*}=1\mid L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$\text{by }Y^{*}\perp\Delta_{A}\mid L_{0}\text{ and }Y^{*}\perp A\mid\Delta_{A}=1,L_{0}$
$=$ $\sum_{l_{0}}\mathbb{P}(Y^{*}=1\mid\Delta_{A}=1,A=a,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$\text{by }Y^{*}\perp\Delta_{Y}\mid L_{1},A=a,\Delta_{A}=1,L_{0}$
$=$ $\sum_{l_{0}}\sum_{l_{1}}\mathbb{P}(Y^{*}=1\mid\Delta_{Y}=1,L_{1}=l_{1},\Delta_{A}=1,A=a,L_{0}=l_{0})\times$
$\quad\mathbb{P}(L_{1}=l_{1}\mid\Delta_{a}=1,A=a,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$=$ $\mathbb{E}\bigg\{\mathbb{E}\big[\mathbb{E}(Y\mid\Delta_{Y}=1,L_{1},A=a,\Delta_{A}=1,L_{0})\mid A=a,\Delta_{A}=1,L_{0})\big]\bigg\}$
where the inner expectation averages out the outcome $Y$ given the conditioning set, the middle expectation average out the time-varying covariates $L_{1}$ given the conditioning set, and the outer expectation averages out the baseline covariates $L_{0}$. For the corresponding statistical estimand to be well-defined, we also need the following positivity assumptions: $\mathbb{P}(\Delta_{Y}=1\mid L_{1},A=a,\Delta_{A}=1,L_{0})>0$ a.e. in addition to the positivity assumptions from the previous section.
Appendix S2.3: Missing Exposures and Outcomes at Baseline and Follow-up (Figure 3 in the main text)
Let $Y_{0}^{*}=Y_{0}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1}$ and $Y_{1}^{*}=Y_{1}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1,\Delta_{Y_{1}}=1}$. Recall that we defined the target parameter for this section as
$\mathbb{P}(Y_{1}^{*}=1\mid Y_{0}^{*}=0)=\frac{\mathbb{P}(Y_{1}^{*}=1,Y_{0}^{*}=0)}{\mathbb{P}(Y_{0}^{*}=0)}$
Using the form of the target parameter on the right-hand side of the above equation, we proceed by presenting a separate identification result for the numerator and denominator separately.
Identification proof for the denominator
Under the following assumptions, which are analogous to Appendix S2.1, we can identify 1 minus the denominator:
$\mathbb{P}(Y_{0}^{*}=1)$ $=$ $\sum_{l_{0}}\mathbb{P}(Y_{0}^{*}=1\mid L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$\text{by }Y_{0}^{*}\perp\Delta_{A}\mid L_{0}\text{ and }Y_{0}^{*}\perp A\mid\Delta_{A}=1,L_{0}\text{ and }Y_{0}^{*}\perp\Delta_{Y_{0}}\mid A=a,\Delta_{a}=1,L_{0}$
$=$ $\sum_{l_{0}}\mathbb{P}(Y_{0}^{*}=1\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$=$ $\mathbb{E}\big[\mathbb{E}(Y\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\big]$
along with the corresponding positivity assumptions.
Identification proof for the numerator
Let $Z^{*}=\mathbb{I}(Y_{1}^{*}=1,Y_{0}^{*}=0)$. Then under the following assumptions, we can identify the numerator $\mathbb{P}(Y_{1}^{*}=1,Y_{0}^{*}=0)=\mathbb{P}(Z^{*}=1)$.
$\mathbb{P}(Z^{*}=1)$ $=$ $\sum_{l_{0}}\mathbb{P}(Z^{*}=1\mid L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$\text{by }Z^{*}\perp\Delta_{A}\mid L_{0}\text{ and }Z^{*}\perp A\mid\Delta_{A}=1,L_{0}\text{ and }Z^{*}\perp\Delta_{Y_{0}}\mid A=a,\Delta_{A}=1,L_{0}$
$=$ $\sum_{l_{0}}\mathbb{P}(Z^{*}=1\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$=$ $\sum_{l_{0}}\sum_{y_{0}}\sum_{l_{1}}\mathbb{P}(Z^{*}=1\mid L_{1}=l_{1},Y_{0}=y_{0},\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\times$
$\quad\mathbb{P}(L_{1}=l_{1},Y_{0}=y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$\text{by }Z^{*}=0\text{ when }Y_{0}=1$
$=$ $\sum_{l_{0}}\sum_{l_{1}}\mathbb{P}(Z^{*}=1\mid L_{1}=l_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\times$
$\quad\mathbb{P}(L_{1}=l_{1},Y_{0}=0\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$\text{by }Z^{*}\perp\Delta_{Y_{1}}\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}$
$=$ $\sum_{l_{0}}\sum_{l_{1}}\mathbb{P}(Z^{*}=1\mid\Delta_{Y_{1}}=1,L_{1}=l_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\times$
$\quad\mathbb{P}(L_{1}=l_{1},Y_{0}=0\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$=$ $\sum_{l_{0}}\sum_{l_{1}}\mathbb{P}(Z^{*}=1\mid\Delta_{Y_{1}}=1,L_{1}=l_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\times$
$\quad\mathbb{P}(L_{1}=l_{1}\mid Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\times$
$\quad\mathbb{P}(Y_{0}=0\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$=$ $\mathbb{E}[\mathbb{E}\left(\mathbb{E}(Y_{1}\mid\Delta_{Y_{1}}=1,L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}\right)]$
For the corresponding statistical estimand to be well-defined, we also need the following positivity assumptions: $P(\Delta_{Y_{1}}=1\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})>0$ a.e. in addition to the positivity assumptions for the denominator.
Appendix S3: Accounting for Outcome Dependence
Here, we outline an approach to account for the dependence of participant outcomes within groups or clusters, such as households, schools, hospitals, or communities. Such dependence could arise due to shared exposures and/or the spread of social behaviors or infectious diseases. In our running example, TB is transmitted from person to person. This dependence should be reflected in the corresponding causal model (e.g., [67, 68, 69, 70]). By following the Causal Roadmap or a similar framework for causal inference [21, 71], we specify causal models encoding our knowledge about the hierarchical data generating process without imposing parametric modeling assumptions — in contrast to more traditional approaches, such as generalizing estimating equations or mixed effects models (e.g., [51, 70, 25, 26, 72]).
Suppose it is reasonable to assume that participant outcomes are dependent within households, but effectively independent between households. Then our causal model would be specified at the household-level, and identification would consider the influence of other household members as well as community-level factors. (See, for example, [70].) Concretely, this may involve including community indicators in $L_{0}$ and summary measures of household-level covariates in $L_{0}$ and $L_{1}$. The exact form of the causal model and identification result will depend on the application. Going forward, we use “cluster” to refer to any group considered to be the (conditionally) independent unit [51, 26, 72].
If clustering is present, estimation and inference must be adjusted. First, the cross-validation scheme used within Super Learner must respect the independent unit. Concretely, participants in a given cluster are all assigned to the same sample-split. Second, for influence curve-based inference, let $m=\{1,\ldots,M\}$ index the clusters and $j=\{1,\ldots,Z_{m}\}$ index for participants in cluster $m$ [18]. Then the total number of participants is $N=\sum_{m}Z_{m}$, and the asymptotic linearity result is re-expressed as $\hat{\Psi}-\Psi=\frac{1}{M}\sum_{m=1}^{M}\left(\sum_{j\in Z_{m}}D_{m,j}\frac{M}{N}\right)$ where $D_{m,j}$ denotes the influence curve for the $j^{th}$ participant in the $m^{th}$ cluster and where we suppressed the second-order remainder term for notational convenience. Altogether, $X_{m}=\frac{M}{N}\sum_{j\in Z_{m}}D_{m,j}$ is the cluster-level influence curve, which has aggregated the individual-level influence curves within cluster $m$ and is weighted by the ratio of the number of clusters to the number of individuals $M/N$. We then proceed with variance estimation using the cluster-level influence curve. This approach is equivalent to using an independent working correlation matrix when obtaining robust (sandwich-based) inference.