### v3

# Causal Inference with Missing Exposures and Missing Outcomes

Abstract
Missing data are ubiquitous in public health research. When estimating causal effects, there are well-established methods to address bias to due missing outcomes. Commonly, causal estimands are defined under hypothetical interventions to “set” the exposure and to prevent missingness. We demonstrate how this framework can be extended to missing exposures. We further extend this framework to incorporate missingness on the baseline outcome, which induces missingness on the population of interest. To do so, we highlight the use of Counterfactual Strata Effects: causal estimands where the focus population is subject to missingness and/or impacted by the exposure. Our work is motivated by SEARCH-TB’s investigation of the effect of alcohol consumption on the risk of incident tuberculosis (TB) infection in rural Uganda. This study posed several real-world challenges: confounding, missingness on the exposure (alcohol use), missingness on the baseline outcome (defining who was at-risk of TB and, thus, in the focus population), and missingness on the outcome at follow-up (capturing who acquired TB). We present a series of causal models and identification results to demonstrate the handling of missingness in these settings. We highlight the use of TMLE with Super Learner and the real-world consequences of our approach.
Keywords: Causal Inference, Counterfactual Strata Effects, Missing Data, Real-world Evidence, Super Learner, Targeted Minimum Loss-based Estimation, TMLE, Tuberculosis
Introduction
Missing data affect the integrity of analyses across the spectrum of public health research, including surveillance studies to estimate disease prevalence and randomized trials to establish efficacy of new medical products.[37, 80, 69, 46, 5, 12, 33, 42] There is rich history of methods research to address the potential for bias when participants with measured outcomes differ meaningfully from those with missing or censored outcomes.[61, 52, 76, 28, 31, 67, 73, 59, 70, 68, 77, 84] The Causal Roadmap provides one such approach.[53, 15, 14, 21] The standard implementation of the Roadmap with censored outcomes is as follows. First, we specify our causal question: how would expected outcomes differ if all did versus did not have the exposure of interest and censoring were prevented. Then, we specify a causal model, such as a directed acyclic graph (DAG) or non-parametric structural equation model (NPSEM), to represent the data generating process for the baseline and time-varying confounders, exposures, and outcomes.[52, 76] Third, we intervene on the causal model to generate counterfactual outcomes under hypothetical interventions to set the exposures and ensure outcome measurement. Fourth, we evaluate whether the corresponding causal estimand can be identified (i.e., if sequential exchangeability and positivity hold).[61, 52, 76, 28] Fifth, we specify the statistical estimand, which is often a complex function of the observed data distribution (i.e., not equal to single regression coefficient). Sixth, we conduct estimation and inference with inverse weighting, standardization, or doubly robust methods, such as targeted minimum loss-based estimation (TMLE).[76] Finally, we conduct sensitivity analyses and appropriately interpret the results.
Here, we first discuss how this framework for missing and censored outcomes can be extended to address missingness on the exposure. We then introduce Counterfactual Strata Effects: causal estimands defined for a group which is subject to missingness and/or influenced by the exposure.[7, 54, 2, 25, 5, 4, 50, 40, 55, 22, 47] Specifically, we highlight the consequences of missingness on the baseline outcome when it is crucial to defining the population of interest. Concretely, suppose we are interested in studying the incidence of some disease. Our focus population would be persons who are at-risk of the developing the outcome and are, thereby, disease-free at baseline. In this setting, our incidence estimates would be subject to bias if our analyses did not account for differential missingness of outcomes at baseline. We provide a framework for explicitly defining, identifying, and estimating parameters in such scenarios—overall and, importantly, when the exposure influences the baseline risk of the outcome. We also discuss how our approach avoids “failures” sometimes arising in the target trial emulation framework.[26]
We illustrate practical relevance with SEARCH-TB, which aimed to evaluate the effect of alcohol use on incident tuberculosis (TB) infection.[1] SEARCH-TB was a sub-study of SEARCH, a cluster randomized trial to evaluate a community-based approach to Universal HIV Test-and-Treat in rural Kenya and Uganda (2013-2017; NCT01864603).[25] At baseline in SEARCH, we conducted a rapid census and then population-based measurement of sociodemographic factors (e.g., age, sex, mobility, and alcohol use) and HIV.[11] Due to high costs and complex logistics, evaluation of incident TB infection was limited to eastern Ugandan and intentionally enriched for persons with HIV.[39, 40] Specifically, we over-sampled households with at least one adult with HIV. Then we aimed to administer tuberculin skin tests to residents of the sampled households. One year later, we aimed to administer follow-up tests to participants who tested negative at baseline in the sub-study. Here, we demonstrate the methods used to address several real-world challenges arising in SEARCH-TB: confounding, missingness on the exposure of interest (self-report of any alcohol use), missingness on the baseline outcome (defining who was at-risk of TB), and missingness on the final outcome (defining who acquired TB).
Building to Counterfactual Strata Effects
Many studies feature only a subset of the challenges described above. We provide causal models and identification results for a series of hypothetical studies with increasing complexity. For simplicity, we focus on defining and identifying causal estimands under a single level of the exposure, but our results naturally generalize to causal effects defined in terms of contrasts of counterfactual outcome distributions under two levels of the exposure (i.e., the average treatment effect or causal risk ratio). In Appendix S1, we review the classic “point-treatment” problem, where we have measured confounding by baseline covariates $L$, a binary exposure $A$ occurring at single time-point, and an outcome $Y$ occurring at the study’s close. Appendix S2 provides a brief review of approaches to handle missingness on the baseline confounders $L$.
Missing Outcomes
Consider our motivating study of the effect of alcohol use on incident TB infection. Suppose that we have a representative cohort of persons without TB at baseline, but some participants did not test for TB at the end of follow-up. Let $\Delta_{Y}$ be an indicator of outcome measurement. If $\Delta_{Y}=1$ for a participant, we observe their outcome $Y$ as usual. However, if $\Delta_{Y}=0$, their outcome $Y$ is not observed. Figure 1 provides two possible causal models to represent such a study. In Panel A, we use $\Upsilon$ to represent the underlying value of the outcome, and define the observed outcome as the product of its measurement indicator and underlying value: $Y=\Delta_{Y}\times\Upsilon$. Thus, $Y$ equals its underlying value $\Upsilon$ when it’s measured (i.e., when $\Delta_{Y}=1$) and is zero otherwise (i.e., when $\Delta_{Y}=0$). In Panel B, we omit the underlying outcome and directly represent the causal model in terms of the observed data: $O=(L,A,\Delta_{Y},Y)$. In the latter presentation, the measurement indicator $\Delta_{Y}$ again influences the value of the observed outcome $Y$, because the outcome is missing if $\Delta_{Y}=0$. Both approaches lead us to same identification assumptions and statistical estimands. Therefore, to minimize notation and mirror the long-standing literature on censoring, we use the latter representation in the remainder of the article.[60, 59, 8, 70, 76, 68, 77, 84, 28]
To define causal effects when the outcome is subject to missingness, we consider counterfactuals indexed by both the exposure and the outcome measurement indicator. Specifically, let $Y^{*}=Y^{A=a,\Delta_{Y}=1}$ be the counterfactual outcome for a given participant if, possibly contrary-to-fact, their exposure were at level $A=a$ and missingness on the outcome prevented. Then our causal estimand $\mathbb{E}(Y^{*})$ is the expected counterfactual outcome if all participants had exposure $A=a$ and their outcomes measured. To identify this causal estimand and express it as a function of the observed data distribution, we would need the baseline covariates $L$ to be sufficient to control for confounding and differential outcome missingness. Specifically, we need $L$ to capture all the common causes of the exposure and outcome, and, among those with the exposure of interest, all the common causes of the outcome and its measurement. These assumptions can be represented as $Y^{*}\perp A\mid L$ and $Y^{*}\perp\Delta_{Y}\mid A=a,L$, respectively. These conditions are often referred to as the “sequential randomization assumption” or “sequential exchangeability” and can be evaluated graphically through the sequential backdoor criterion.[61, 52, 76, 28] We also need a positive probability of being exposed to level $A=a$ within all possible values of the confounders and, among those with the exposure of interest, a positive probability of outcome measurement within all possible values of the confounders: $\mathbb{P}(A=a\mid L)>0$ a.e. and $\mathbb{P}(\Delta_{Y}\mid A=a,L)>0$ a.e., respectively. If these assumptions hold, we have equivalence between our causal estimand $\mathbb{E}(Y^{*})$ and the statistical estimand given by the G-computation formula: $\mathbb{E}[\mathbb{E}(Y\mid\Delta_{Y}=1,A=a,L)]$.[61]
Figure 1: Causal models with missingness on the outcome (Panels A-B) or missingness on the exposure (Panel C): LL=baseline covariates, $\Delta_{A}$=indicator of exposure measurement, AA=observed exposure, $\Upsilon$=underlying outcome, $\Delta_{Y}$=indicator of outcome measurement, and YY= observed outcome. For ease of presentation, the models are shown without dependence between the unmeasured variables, which are omitted on the graph.
Missing Exposures
We now demonstrate how the above approach for missing outcomes can be applied to missing exposures. Continuing our running example, suppose that among our representative cohort of persons without TB at baseline, some participants did not answer questions about their alcohol use. Let $\Delta_{A}$ be an indicator that a participant has their exposure measured. If $\Delta_{A}=1$ for a participant, we observe their exposure $A$ as usual. However, if $\Delta_{A}=0$ for a participant, their exposure $A$ is missing. The DAG and NPSEM for such a study are given in Figure 1C. To define causal effects when the exposure is subject to missingness, we consider counterfactuals indexed by both the exposure and its measurement indicator. Now, let $Y^{*}=Y^{\Delta_{A}=1,A=a}$ be the counterfactual outcome for a given participant if, possibly contrary-to-fact, their exposure were measured and at level $A=a$.
Then to identify the expected counterfactual outcome $\mathbb{E}(Y^{*})$ and express it as a function of the distribution of the observed data $O=(L,\Delta_{A},A,Y)$, we need analogous conditions as the prior subsection. Specifically, we need $L$ to be sufficient to control for differential exposure missingness and for confounding (among those with measured exposures): $Y^{*}\perp\Delta_{A}\mid L$ and $Y^{*}\perp A\mid\Delta_{A}=1,L$, respectively. We also need the two analogous assumptions on data support: $\mathbb{P}(\Delta_{A}=1\mid L)>0$ a.e. and $\mathbb{P}(A=a\mid\Delta_{A}=1,L)>0$ a.e. If these four assumptions hold, we can rewrite $\mathbb{E}(Y^{*})$ as the statistical estimand $\mathbb{E}[\mathbb{E}(Y\mid A=a,\Delta_{A}=1,L)]$ with proof in Appendix S3.1.
Missing Exposures and Outcomes
We now combine our two challenges: missing exposures and missing outcomes. Continuing our running example, suppose that among our cohort of persons at-risk of TB, some participants did not answer questions about their alcohol use and some participants did not test for TB at the end of follow-up. To reflect this data generating process, we introduce new notation to reflect the longitudinal setting. Let $L_{0}$ be baseline covariates and $L_{1}$ be additional covariates collected after the exposure but before outcome ascertainment. Figure 2 provides the causal models for such a study.
Figure 2: Causal graph and corresponding non-parametric structural equation model with missingness on the exposure and the outcome: $L_{0}$=baseline covariates, $\Delta_{A}$=indicator of exposure measurement, AA=observed exposure, $L_{1}$=time-varying covariates, $\Delta_{Y}$=indicator of outcome measurement, and YY=observed outcome. For ease of presentation, the models are shown without dependence between the unmeasured variables, which are omitted on the graph.
To define the causal effect when the exposure and outcome are subject to missingness, we now consider counterfactuals indexed by the exposure and two measurement indicators. Specifically, let $Y^{*}=Y^{\Delta_{A}=1,A=a,\Delta_{Y}=1}$ denote the counterfactual outcome under hypothetical interventions to ensure exposure measurement, “set” the exposure level to $A=a$, and ensure outcome measurement. To identify the expected counterfactual outcome $\mathbb{E}(Y^{*})$ and express it as a function of the distribution of the observed data $O=(L_{0},\Delta_{A},A,L_{1},\Delta_{Y},Y)$, we now need to account for the post-baseline covariates $L_{1}$, which act as time-dependent confounders. Specifically, $L_{1}$ are mediators of the exposure-outcome relationship, while “confounding” the measurement-outcome relationship. Therefore, we rely on sequential randomization/exchangeability and find a set of covariates that satisfies the backdoor criterion for each “intervention” node given the observed past.[61] As before, we need that the baseline covariates $L_{0}$ are sufficient to control for missing exposures and for confounding. In other words, we need the analogous identification assumptions given in the prior subsection. Additionally, we need that among participants with measured exposures at the level of interest (i.e., $\Delta_{A}=1$ and $A=a$), the baseline and time-varying covariates ($L_{0},L_{1})$ capture all the common causes of outcomes and their measurement as well as a positive probability of outcome measurement within all possible values of the baseline and time-varying covariates: $Y^{*}\perp\Delta_{Y}\mid L_{1},A=a,\Delta_{A}=1,L_{0}$ and $\mathbb{P}(\Delta_{Y}=1\mid L_{1},A=a,\Delta_{A}=1,L_{0})>0$ a.e., respectively. If these assumptions hold, we can rewrite $\mathbb{E}(Y^{*})$ in terms of the longitudinal G-computation formula: $\mathbb{E}\{\mathbb{E}[\mathbb{E}(Y\mid\Delta_{Y}=1,L_{1},A=a,\Delta_{A}=1,L_{0})\mid A=a,\Delta_{A}=1,L_{0})]\}$, shown in terms of iterated expectations and with proof in Appendix S3.2.[61, 8, 72]
Counterfactual Strata Effects
Following our motivating example, we now introduce missingness on the outcome at baseline, which induces missingness on the focus population. Specifically in SEARCH-TB, we were interested in the effect of alcohol use among all persons at-risk of TB, but we did not reach all for initial TB testing. Therefore, our longitudinal cohort of participants who tested TB-negative at baseline in the sub-study was not representative of all persons at-risk of TB. Furthermore, alcohol use likely influenced baseline TB status.[38, 58, 56] Under the target trial framework, SEARCH-TB would result in “Emulation failure 2: time zero is set at eligibility but after strategy assignment”.[26] However, we now demonstrate how following the Causal Roadmap facilitates clear definition of the causal estimand with Counterfactual Strata Effects, while avoiding bias due to missing data during estimation and inference.
To do so, we update our notation to have multiple outcome measures. Let $\Delta_{Y_{0}}$ be an indicator of outcome measurement at start of follow-up (hereafter “baseline”) and $\Delta_{Y_{1}}$ be an indicator of outcome measurement at the end of follow-up (hereafter “endline”). Let $Y_{0}$ and $Y_{1}$ denote the corresponding outcomes. The corresponding causal models can be found in Figure 3. The time-ordering reflects the study protocol and procedures in SEARCH-TB; specifically, alcohol use $A$ was measured before baseline TB status $Y_{0}$.[39, 40, 1] As a result, the exposure $A$ can impact who has prevalent TB at baseline ($Y_{0}=1$) and, thereby, who is at-risk of TB at baseline ($Y_{0}=0$).
Figure 3: Causal graph and corresponding non-parametric structural equation model with missingness on the exposure and the outcome at baseline and at endline: $L_{0}$=baseline covariates, $\Delta_{A}$=indicator of exposure measurement, A=exposure, $\Delta_{Y_{0}}$=indicator of outcome measurement at baseline, $Y_{0}$=baseline outcome, $L_{1}$=time-dependent covariates, $\Delta_{Y_{1}}$=indicator of outcome measurement at endline, and $Y_{1}$= outcome at endline. For ease of presentation, the models are shown without dependence between the unmeasured variables, which are omitted on the graph.
To define the causal estimand in this setting, we first consider the counterfactual outcome at baseline under hypothetical interventions to ensure exposure measurement, “set” the exposure level to $A=a$, and ensure outcome measurement at baseline: $Y_{0}^{*}=Y_{0}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1}$. Additionally, we consider the counterfactual outcome at endline under the prior interventions as well as a hypothetical intervention to ensure outcome measurement at endline among those known to be at-risk at baseline. Concretely, only participants who tested TB-negative at baseline ($\Delta_{Y_{0}}=1$ and $Y_{0}=0$) were approached for re-testing at endline. Thereby, this final intervention is a dynamic or personalized.[7, 29, 74, 62] We set $\Delta_{Y_{1}}$ equal to one if $Y_{0}=0$ and to zero otherwise. For simplicity, denote the resulting counterfactual outcome as $Y_{1}^{*}=Y_{1}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1,\Delta_{Y_{1}}=1}$.
Now we can precisely define the causal estimand in terms of the following conditional probability, which captures the counterfactual incidence of the outcome among all at-risk at baseline: $\mathbb{P}(Y_{1}^{*}=1\mid Y_{0}^{*}=0)$. Due to conditioning on a counterfactual variable, such parameters are sometimes called “Counterfactual Strata Effects”, which are defined by Nakato et al. as “causal estimands where the outcome is only relevant for a group whose membership is subject to missingness and/or impacted by the exposure”.[47] These estimands have previously been specified to evaluate intervention effects in cluster randomized trials.[7, 54, 2, 25, 5, 4, 50, 40, 55, 22, 47] To our knowledge, this is their first use in observational studies and with additional missingness on the exposure.
To facilitate identification, we re-express the conditional probability as
$\mathbb{P}(Y_{1}^{*}=1\mid Y_{0}^{*}=0)=\frac{\mathbb{P}(Y_{1}^{*}=1,Y_{0}^{*}=0)}{\mathbb{P}(Y_{0}^{*}=0)}$ (1)
Then given the observed data $O=(L_{0},\Delta_{A},A,\Delta_{Y_{0}},Y_{0},L_{1},\Delta_{Y_{1}},Y_{1})$, we can identify the denominator and the numerator. The denominator $\mathbb{P}(Y_{0}^{*}=0)$ represents the counterfactual prevalence of not having TB at baseline — under alcohol use $A=a$ and under complete measurement of the exposure and baseline outcome ($\Delta_{A}=\Delta_{Y_{0}}=1)$. The causal structure for this parameter is analogous to that of Section 3.2, but with an additional measurement indicator for the baseline outcome. Therefore, under analogous assumptions, we can identify $\mathbb{E}(Y_{0}^{*})=\mathbb{E}\big[\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\big]$ with proof in Appendix S3.3. Since we are interested the counterfactual probability of being at-risk at baseline $\mathbb{P}(Y_{0}^{*}=0)$, our statistical estimand for the denominator becomes $1-\mathbb{E}\big[\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\big]$. Altogether, this statistical estimand accounts for differential measurement of the exposure and baseline TB status as well as the impact of alcohol use $A=a$ on baseline TB risk.
In our final causal parameter (Eq. 1), the numerator $\mathbb{P}(Y_{1}^{*}=1,Y_{0}^{*}=0)$ represents the counterfactual probability of acquiring TB — under alcohol use $A=a$ and under complete measurement $\Delta_{A}=\Delta_{Y_{0}}=\Delta_{Y_{1}}=1$. For ease of notation, let $Z^{*}=\mathbb{I}(Y_{1}^{*}=1,Y_{0}^{*}=0)$ represent the joint indicator of these two counterfactual values. To identify $\mathbb{E}(Z^{*})=\mathbb{P}(Z^{*}=1)$, we need analogous assumptions as for the denominator together with the following. Among those known to be at-risk at baseline ($\Delta_{Y_{0}}=1,Y_{0}=0$) and with measured exposure of interest ($\Delta_{A}=1,A=a$): the baseline and time-varying covariates capture the common causes of the joint outcome and endline measurement: $Z^{*}\perp\Delta_{Y_{1}}\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}$. We also need there to be a positive probability of follow-up measurement within all possible values of $L_{0}$ and $L_{1}$: $\mathbb{P}(\Delta_{Y_{1}}=1\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})>0$ a.e. Under these assumptions and with proof given in Appendix S3.3, the numerator is identified as
$\mathbb{P}(Z^{*}=1)\ =\ \mathbb{E}[\mathbb{E}\left(\mathbb{E}(Y_{1}\mid\Delta_{Y_{1}}=1,L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}\right)]$
Putting it all together, our statistical estimand is given by
$\Psi(\mathbb{P};a)=\frac{\mathbb{E}[\mathbb{E}\left(\mathbb{E}(Y_{1}\mid\Delta_{Y_{1}}=1,L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}\right)]}{1-\mathbb{E}\left[\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\right]}$ (2)
for exposure level $A=a$. Then we can define associations in terms of contrasts $\Psi(\mathbb{P};a)$ at different exposure levels. Concretely, in SEARCH-TB, we were interested in evaluating the association of alcohol consumption on incident TB infection with the risk ratio: $\Psi(\mathbb{P})=\Psi(\mathbb{P};1)\div\Psi(\mathbb{P};0)$.
Statistical Estimation and Inference
In the previous section, we introduced a series of causal models and identification results of increasing complexity. For the resulting statistical estimands, we could use singly robust estimators, such as standardization (a.k.a., “G-computation”) or inverse probability weighting (IPW).[61, 31] Here, we highlight the use of TMLE, which is a doubly robust procedure and asymptotically efficient under certain conditions.[76] In TMLE, initial estimates of the relevant pieces of the observed data distribution are updated to achieve the optimal bias-variance trade-off for the estimand and to solve the efficient influence curve equation. Initial estimates are often computed via Super Learner, an ensemble machine learning algorithm using V-fold cross-validation to select an optimal weighted linear combination of predictions from a library of candidate learners.[75] Thereby, TMLE leverages machine learning to avoid introducing new modeling assumptions during estimation, while supporting valid statistical inference under reasonable conditions. Notably, for ratio-type estimands corresponding to Counterfactual Strata Effects (Eq. Appendix S4.1: Simulation Results), we would implement a separate TMLE for the numerator estimand (the joint probability) and the denominator estimand (1- the baseline prevalence) before combining the results.
TMLE is an asymptotically linear estimator and is normally distributed in the large data limit.[76] The estimator minus the estimand behaves like a sample mean in the first order: $\hat{\Psi}-\Psi=\frac{1}{N}\sum_{i=1}^{N}D_{i}+o_{P}(N^{-1/2})$ where $D_{i}$ is the influence curve for participant $i=\{1,\ldots,N\}$ and $o_{P}(N^{-1/2})$ is a second-order remainder term going to zero in probability.[79] The estimated influence curve is used to calculate standard errors, Wald-type confidence intervals, and p-values. Concretely, a 95% confidence interval is constructed using $\hat{\Psi}\pm z_{0.0975}\frac{\hat{\sigma}}{\sqrt{n}}$ where $z_{0.975}$ is the critical value at the 97.5th-percentile of the standard normal and $\hat{\sigma}$ is the standard deviation of the estimated influence curve. For ratio-type estimands (Eq. Appendix S4.1: Simulation Results), once the influence curves for the numerator and denominator have been estimated, the Delta method provides an estimate of the influence curve for our overall estimand.[79, 45] Then to calculate measures of association on the difference, ratio, or odds ratio scale, we apply the Delta method a second time to get inference for these types of functionals (Appendix S4).
To evaluate the practical performance of our approach, we conducted a simulation study reflecting the motivating study (Figure 3). The simulation design and results are detailed in Appendix S5. Briefly, we found that TMLE had negligible bias and achieved nominal confidence interval coverage for the exposure-specific estimands $\Psi(\mathbb{P};a)$, given in Eq. Appendix S4.1: Simulation Results for $A=\{0,1\}$, as well as their ratio (i.e. the relative effect) and their difference (i.e., the average treatment effect). All other approaches exhibited meaningful bias and sub-nominal confidence interval coverage.
Application to SEARCH-TB
We now return to our motivating question: what is the effect of alcohol use on incident TB infection among adults in rural Eastern Uganda? With our multinational and interdisciplinary team, we worked through the Causal Roadmap to specify the Statistical Analysis Plan.[53, 15, 14, 21, 48, 1] Our causal model reflected the team’s knowledge of the study protocol, study procedures, and epidemiology of TB in the region. Our adjustment set included the SEARCH trial arm, community indicators, household HIV status, as well as individual-level age, sex, and mobility measures. The exposure was self-report of any alcohol use, and the outcome was incident TB infection: conversion from a negative tuberculin skin test to a positive one.[1] The covariates and exposure were measured at the start of the parent SEARCH study (2013-2014), while TB status was measured subsequently in the sub-study (2015-2017).[25, 39, 40] For the primary analysis, we used TMLE with Super Learner to combine estimates from generalized linear models, multivariate adaptive regression splines, and the mean. We conducted influence curve-based inference, accounting for clustering by household (Appendix S6).[78, 50]
In secondary analyses, we considered communities, instead of households, to be the independent unit. To examine the sensitivity of our results to alternative estimation approaches, we implemented two singly robust approaches for the final statistical estimand (Eq. Appendix S4.1: Simulation Results): G-Computation using parametric regressions to estimate the iterated conditional expectations and IPW using parametric regressions to estimate the weights.[8] Finally, to examine the impact of our handling of missing data, we took the following “naïve” approach: (1) subset on participants with complete data $\Delta_{A}=\Delta_{Y_{0}}=\Delta_{Y_{1}}=1$; (2) further subset on participants known to be at-risk at baseline $Y_{0}=0$, and (3) implement TMLE to adjust for confounding and to estimate $\mathbb{E}[\mathbb{E}(Y_{1}\mid A=a,L)]$ among the remaining subset of participants. This approach is inherently flawed, because of the bias induced by inappropriately conditioning on various colliders and mediators (Figure 3). Nonetheless, the approach could be one taken by an analyst aiming to implement a “complete-case” analysis while adjusting for confounding.[64, 41, 17] For statistical inference, G-computation, IPW, and the naïve approach accounted for clustering by household.
In the primary analysis using TMLE with clustering by household, we found that alcohol use was associated with a 49% increase in the risk of incident TB: risk ratio (RR)=1.49 (95%CI: 1.39-1.59).[1] As shown in Figure 4, secondary analyses with the community as the independent unit yielded very similar results, despite meaningfully reducing the effective sample size to 9 communities: RR=1.49 (95%CI: 1.37-1.62); Appendix S6. In contrast, G-computation resulted in a larger association with large confidence intervals, while IPW resulted in a smaller association and confidence intervals overlapping the null: RR=1.58 (95%CI: 1.36-1.83) and RR=1.13 (95%CI: 1.00-1.27), respectively. Finally, after restricting to participants who responded to questions about alcohol use, tested negative at baseline, and tested again at follow-up, the naïve approach was the least precise and resulted in the widest confidence intervals: RR=1.18 (95%CI: 0.89-1.57).
Figure 4: Results from SEARCH-TB for the association of alcohol use on incident tuberculosis (TB) infection: “Primary” with TMLE and clustering by household, “Secondary” with TMLE and clustering by community, “G-Comp” with G-computation, “IPW” with inverse probability weighting, and “Naïve” based on subsetting on those at-risk at baseline and with measured exposures and outcomes at endline.
Discussion
We presented causal models, causal estimands, and identification results for a series of prospective studies with increasing levels of missingness. We reviewed how the Causal Roadmap could be applied to address missing outcomes and extended to address missing exposures. We then motivated the use of Counterfactual Strata Effects to precisely define causal estimands when the focus population is subject to missingness and/or impacted by the exposure. Specifically, we considered the common scenario where there is missingness on the outcome at baseline — determining who is at-risk of acquiring the outcome by endline. We further allowed for the exposure to influence baseline outcome risk. We demonstrated that while these real-world complications would result in “failures” in the target trial emulation framework,[26] our Roadmap approach facilitates definition and identification of the corresponding causal estimand. For estimation and inference, we highlighted the use of TMLE with Super Learner to robustly and efficiently estimate the corresponding statistical estimands.
We emphasize that Counterfactual Strata Effects are distinct from Principal Strata Effects.[16, 18, 19, 51] For simplicity, suppose that we did not have missingness, and define $Y_{0}^{a}$ and $Y_{1}^{a}$ as the counterfactual outcome at baseline and endline under exposure $A=a$. Then in SEARCH-TB, principal stratification could be applied to define the effect of alcohol on incident TB among the unobservable subset of participants who would have always been at-risk of TB at baseline regardless of their alcohol use: $\mathbb{P}(Y_{1}^{1}=1\mid Y_{0}^{1}=Y_{0}^{0}=0)$ versus $\mathbb{P}(Y_{1}^{0}=0\mid Y_{0}^{1}=Y_{0}^{0}=0)$. Instead, Counterfactual Strata Effects enable us to define the effect of alcohol on incident TB among the counterfactual population of persons at-risk at baseline under alcohol use $A=a$: $\mathbb{P}(Y_{1}^{1}=1\mid Y_{0}^{1}=0)$ versus $\mathbb{P}(Y_{1}^{0}=1\mid Y_{0}^{0}=0)$. The at-risk population is exposure-specific, and our estimand captures the effect of alcohol use on prevalent TB at baseline and on incident TB by endline. Of course, we could also directly examine the effect on prevalent TB at baseline: $\mathbb{P}(Y_{0}^{1}=1)$ versus $\mathbb{P}(Y_{0}^{0}=1)$. In addition to incidence studies,[40, 50] Counterfactual Strata Effects commonly arise when studying care cascades: the series of steps from screening, through diagnosis and treatment, to control.[7, 54, 2, 25, 5, 4, 47] Indeed, these cascades are a series of conditional probabilities — each subject to missingness and each potentially impacted by the exposure or intervention strategy.
There are several advantages to our framework for handling multi-source missingness. As previously discussed, our approach avoids a common “failure” arising when aiming to emulate at target trial.[26] Furthermore, unlike the m-DAG approach which aids in identifying the entire joint distribution of the observed data,[46] our approach focuses our efforts on the portion of the observed data distribution relevant for the statistical estimand (Appendix S7). Our approach also uses data on all participants, improving efficiency relative to approaches excluding participants with missing data on the relevant variables (Figure 4; Appendix S.5). Finally, our approach leads us to statistical estimands that can be robustly and rigorously estimated with modern methods, such as TMLE.
There are several limitations to our work. First, we focused on cross-sectional or prospective studies. Thus, we did not cover scenarios where the outcome impacts the measurement of other variables. Such scenarios would arise in case-control studies and have been addressed in prior literature.[65, 63, 86, 34, 17, 57] Second, we did not provide an exhaustive set of causal models and identification results for all possible studies; however, our approach is generalizable and covers many scenarios arising in public health. Third, we did not consider multiple imputation, which is a common approach for missing data.[66] However, foundational work is needed to investigate the assumptions, implementation, and performance of multiple imputation in settings mirroring our motivating example: (1) missingness on the exposure, (2) missingness on the baseline outcome, (3) missingness in the final outcome, (4) confounding, and (5) dependence among study participants. Indeed, Tompsett et al. suggested that multiple imputation cannot be used in settings with missingness on the focus population (termed “missing eligibility data”) and missingness on the exposure.[71] Finally, we relied on various versions of the sequential exchangeability assumption for both confounding and missingness. In practice, data may be missing as a result of unobserved variables, and we may need to collect additional data as well as conduct sensitivity/bias analyses.[36, 13] Nonetheless, even when a “causal gap” remains, we have a framework to define a statistical estimand, which is aligned with our research question, and appropriately addresses many real-world complications.[53, 15, 14, 21, 48]
Supplementary Materials for “Causal Inference with Missing Exposures and Missing Outcomes”
Contents:
  • Appendix S1: The classic point-treatment problem
  • Appendix S2: Missing confounders
  • Appendix S3: Proofs
  • Appendix S4: More on the Delta method
  • Appendix S5: Finite sample simulation study
  • Appendix S6: Accounting for outcome dependence
  • Appendix S7: Missingness DAGs
Appendix S1: The classic point-treatment problem
We consider the classic “point-treatment” problem, where we have measured confounding by baseline covariates $L$, a binary exposure $A$ occurring at single time-point, and an outcome $Y$ occurring at the study’s close. This could represent a study of the effect of alcohol use ($A$) on incident TB infection ($Y$) among a representative cohort of persons without TB at baseline. The directed acyclic graph (DAG) and non-parametric structural equation model (NPSEM) for such a study are given in Figure 5A.
Under interventions on the causal model, we generate counterfactual outcomes corresponding to the research question of interest. Specifically, let $Y^{a}$ be the counterfactual outcome for a given participant if, possibly contrary-to-fact, they had exposure-level $A=a$. Then our causal estimand $\mathbb{E}(Y^{a})$ is the expected counterfactual outcome if all had exposure-level $A=a$. In our running example, $\mathbb{E}(Y^{a})=\mathbb{P}(Y^{a}=1)$ is the counterfactual risk of incident TB infection with alcohol use $A=a$.
Figure 5: Causal models for a classic point-treatment problem with complete measurement of the baseline covariates $L$, the exposure $A$, and the outcome $Y$. On the directed acyclic graph, $U^{*}$ represents unmeasured common causes of at least two variables in $(L,A,Y)$. In panel A provides the causal models under no assumptions about the unmeasured factors. Panels B and C are compatible with the no unmeasured confounders assumption.
To identify our causal target parameter and express it as function of the distribution of the observed data $O=(L,A,Y)$, we would need there to be no unmeasured confounding, which corresponds to the assumption that the baseline covariates $L$ capture all the joint causes of the exposure $A$ and outcome $Y$. This condition is called “the randomization assumption” or “exchangeability”, can be evaluated with the backdoor criterion, and can be represented as $Y^{a}\perp A\mid L$.[61, 52, 76, 28] Concretely, $\mathbb{E}(Y^{a})$ is not identified in Figure 5A because there is an unmeasured common cause, represented by $U^{*}$, of the exposure $A$ and outcome $Y$. In Figure 5B and C, we show two causal models where this assumption would hold.
Additionally, we need there to be a non-zero probability of having the exposure in all possible values of $L$: $\mathbb{P}(A=a\mid L)>0$ a.e. This is a condition on data support and known as the “positivity assumption”. We note that the consistency assumption, stating that the counterfactual outcome $Y^{a}$ equals the observed outcome $Y$ under exposure $A=a$, holds by design when we define counterfactuals through intervention on the causal model. Under the exchangeability and positivity assumptions, our causal target is equal to the statistical estimand given by the G-computation formula: $\mathbb{E}[\mathbb{E}(Y\mid A=a,L)]$.[61] Even if these assumptions are not reasonable (e.g., Figure 5A reflects reality), we still have well-defined statistical estimand, on which we can focus our estimation efforts. In other words, we still proceed with estimation and inference $\mathbb{E}[\mathbb{E}(Y\mid A=a,L)]$ , while appropriately accounting for lack of identification in our interpretation.[53, 15, 21, 82, 14] For ease of presentation, we provide the causal models without dependence between the unmeasured factors in the main text.
Appendix S2: Missing Confounders
Throughout the main text, we also assume that the covariates are completely measured. While missingness on covariates is common, a detailed discussion is beyond the scope of this manuscript, which focuses on defining, identifying, and estimating causal effects in settings with missing exposures and missing outcomes. Here, we provide a brief review of methods to handle missing confounders. Common approaches include dropping covariates subject to missingness, conducting a complete (available) case analysis, including a missingness indicator, and using multiple imputation.[20, 83, 81, 57] Excluding covariates subject to missingness will induce bias if those covariates are confounders. A complete case analysis, where observations with missing data are excluded, is guaranteed to reduce power and will be biased unless missingness is completely at random. In the missing indicator method, a dummy variable is included to indicate the covariate is measured/missing. Recently, the missingness indicator method was shown to perform similarly to multiple imputation (MI) in many settings.[83] MI uses observed variables to predict the values of missing ones and has traditionally relied on strong modeling assumptions, although there is a growing interest in using machine learning.[67, 35, 10]
Williamson et al. provide an excellent discussion and evaluation of more modern methods for missing confounders, including doubly robust approaches of generalized raking and targeted minimum loss-based estimation (TMLE).[81] These approaches are more robust to model misspecification and are efficient in certain scenarios. Williamson et al. also provide guidance on selecting the approach to handle missing confounders based on the bias-variance trade-off.[81] Recently, Qiu et al. extended this work and showed TMLE, a plug-in estimator, had superior performance to generalized ranking in finite samples.[57]
Appendix S3: Proofs
In the following, we provide proofs for the identification results. To match the applied example, we focus on binary outcomes, but our results generalize to all outcome-types. For simplicity we focus on categorical covariates, but our summations generalize to integrals for continuous covariates.
Appendix S2.1: Missing exposures (Figure 1C in the main text)
Let $Y^{*}=Y^{\Delta_{A}=1,A=a}$. Then we have equivalence between our wished-for causal estimand and the corresponding statistical estimand under the following identifability assumptions:
$\mathbb{P}(Y^{*}=1)$ $=$ $\sum_{l}\mathbb{P}(Y^{*}=1\mid L=l)\mathbb{P}(L=l)$
$\text{by }Y^{*}\perp\Delta_{A}\mid L$
$=$ $\sum_{l}\mathbb{P}(Y^{*}=1\mid\Delta_{A}=1,L=l)\mathbb{P}(L=l)$
$\text{by }Y^{*}\perp A\mid\Delta_{A}=1,L$
$=$ $\sum_{l}\mathbb{P}(Y^{*}=1\mid A=a,\Delta_{A}=1,L=l)\mathbb{P}(L=l)$
by the consistency assumption
$=$ $\sum_{l}\mathbb{P}(Y=1\mid A=a,\Delta_{A}=1,L=l)\mathbb{P}(L=l)$
$=$ $\mathbb{E}\big[\mathbb{E}(Y\mid A=a,\Delta_{A}=1,L)\big]$
We again note the consistency assumption holds by our definition of counterfactual outcomes as being derived through interventions on the causal model. For the corresponding statistical estimand to be well-defined, we also need the following positivity assumptions: $\mathbb{P}(\Delta_{A}=1|L)>0$ a.e. and $\mathbb{P}(A\mid\Delta_{A}=1,L)>0$ a.e..
Appendix S2.2: Missing Exposures and Outcomes (Figure 2 in the main text)
Let $Y^{*}=Y^{\Delta_{A}=1,A=a,\Delta_{Y}=1}$. Then we have equivalence between our wished-for causal estimand and the corresponding statistical estimand under the following identifability assumptions:
$\mathbb{P}(Y^{*}=1)$ $=$ $\sum_{l_{0}}\mathbb{P}(Y^{*}=1\mid L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$\text{by }Y^{*}\perp\Delta_{A}\mid L_{0}\text{ and }Y^{*}\perp A\mid\Delta_{A}=1,L_{0}$
$=$ $\sum_{l_{0}}\mathbb{P}(Y^{*}=1\mid A=a,\Delta_{A}=1,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$\text{by }Y^{*}\perp\Delta_{Y}\mid L_{1},A=a,\Delta_{A}=1,L_{0}$
$=$ $\sum_{l_{0}}\sum_{l_{1}}\mathbb{P}(Y^{*}=1\mid\Delta_{Y}=1,L_{1}=l_{1},A=a,\Delta_{A}=1,L_{0}=l_{0})\times$
$\quad\mathbb{P}(L_{1}=l_{1}\mid A=a,\Delta_{A}=1,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$=$ $\mathbb{E}\bigg\{\mathbb{E}\big[\mathbb{E}(Y\mid\Delta_{Y}=1,L_{1},A=a,\Delta_{A}=1,L_{0})\mid A=a,\Delta_{A}=1,L_{0})\big]\bigg\}$
where the inner expectation averages out the outcome $Y$ given the conditioning set, the middle expectation average out the time-varying covariates $L_{1}$ given the conditioning set, and the outer expectation averages out the baseline covariates $L_{0}$. For the corresponding statistical estimand to be well-defined, we also need the following positivity assumptions: $\mathbb{P}(\Delta_{Y}=1\mid L_{1},A=a,\Delta_{A}=1,L_{0})>0$ a.e. in addition to the positivity assumptions from the previous section.
Appendix S2.3: Missing Exposures and Outcomes at Baseline and Follow-up (Figure 3 in the main text)
Let $Y_{0}^{*}=Y_{0}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1}$ and $Y_{1}^{*}=Y_{1}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1,\Delta_{Y_{1}}=1}$. Recall that we defined the target parameter for this section using Counterfactual Strata Effects:[3, 54, 5, 4, 50, 55, 22, 25, 40]
$\mathbb{P}(Y_{1}^{*}=1\mid Y_{0}^{*}=0)=\frac{\mathbb{P}(Y_{1}^{*}=1,Y_{0}^{*}=0)}{\mathbb{P}(Y_{0}^{*}=0)}$
Using the form of the target parameter on the right-hand side of the above equation, we proceed by presenting a separate identification result for the numerator and denominator separately.
Identification proof for the denominator
Under the following assumptions, which are analogous to Appendix S2.1, we can identify 1 minus the denominator:
$\mathbb{P}(Y_{0}^{*}=1)$ $=$ $\sum_{l_{0}}\mathbb{P}(Y_{0}^{*}=1\mid L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$\text{by }Y_{0}^{*}\perp\Delta_{A}\mid L_{0}\text{ and }Y_{0}^{*}\perp A\mid\Delta_{A}=1,L_{0}\text{ and }Y_{0}^{*}\perp\Delta_{Y_{0}}\mid A=a,\Delta_{a}=1,L_{0}$
$=$ $\sum_{l_{0}}\mathbb{P}(Y_{0}^{*}=1\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$=$ $\mathbb{E}\big[\mathbb{E}(Y\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\big]$
along with the corresponding positivity assumptions.
Identification proof for the numerator
Let $Z^{*}=\mathbb{I}(Y_{1}^{*}=1,Y_{0}^{*}=0)$. Then under the following assumptions, we can identify the numerator $\mathbb{P}(Y_{1}^{*}=1,Y_{0}^{*}=0)=\mathbb{P}(Z^{*}=1)$.
$\mathbb{P}(Z^{*}=1)$ $=$ $\sum_{l_{0}}\mathbb{P}(Z^{*}=1\mid L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$\text{by }Z^{*}\perp\Delta_{A}\mid L_{0}\text{ and }Z^{*}\perp A\mid\Delta_{A}=1,L_{0}\text{ and }Z^{*}\perp\Delta_{Y_{0}}\mid A=a,\Delta_{A}=1,L_{0}$
$=$ $\sum_{l_{0}}\mathbb{P}(Z^{*}=1\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$=$ $\sum_{l_{0}}\sum_{y_{0}}\sum_{l_{1}}\mathbb{P}(Z^{*}=1\mid L_{1}=l_{1},Y_{0}=y_{0},\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\times$
$\quad\mathbb{P}(L_{1}=l_{1},Y_{0}=y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$\text{by }Z^{*}=0\text{ when }Y_{0}=1$
$=$ $\sum_{l_{0}}\sum_{l_{1}}\mathbb{P}(Z^{*}=1\mid L_{1}=l_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\times$
$\quad\mathbb{P}(L_{1}=l_{1},Y_{0}=0\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$\text{by }Z^{*}\perp\Delta_{Y_{1}}\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}$
$=$ $\sum_{l_{0}}\sum_{l_{1}}\mathbb{P}(Z^{*}=1\mid\Delta_{Y_{1}}=1,L_{1}=l_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\times$
$\quad\mathbb{P}(L_{1}=l_{1},Y_{0}=0\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$=$ $\sum_{l_{0}}\sum_{l_{1}}\mathbb{P}(Z^{*}=1\mid\Delta_{Y_{1}}=1,L_{1}=l_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\times$
$\quad\mathbb{P}(L_{1}=l_{1}\mid Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\times$
$\quad\mathbb{P}(Y_{0}=0\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
$=$ $\mathbb{E}[\mathbb{E}\left(\mathbb{E}(Y_{1}\mid\Delta_{Y_{1}}=1,L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}\right)]$
For the corresponding statistical estimand to be well-defined, we also need the following positivity assumptions: $P(\Delta_{Y_{1}}=1\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})>0$ a.e. in addition to the positivity assumptions for the denominator.
Appendix S4: The Delta Method for Ratio-Estimands
The Delta method is commonly applied to derive influence curve-based inference for asymptotically linear estimators.[79] Recall our exposure-specific estimand is given by the following ratio:
$\Psi(\mathbb{P};a)=\frac{\mathbb{E}[\mathbb{E}\left(\mathbb{E}(Y_{1}\mid\Delta_{Y_{1}}=1,L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}\right)]}{1-\mathbb{E}\left[\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\right]}$ (3)
Following the notation of Moore and van der Laan,[45] we denote the numerator with $\mu_{1}$ and the denominator with $\mu_{0}$. Let $IC_{1}$ and $IC_{0}$ be the corresponding influence curves. Then based on the Delta method, the influence curve for $log[\Psi(\mathbb{P};a)]$ is given by $IC(a)=1/\mu_{1}\times IC_{1}-1/\mu_{0}\times IC_{0}$. We obtain a variance estimate with the sample variance of $IC(a)$ scaled by sample size $N$. Using that variance estimate, we conduct hypothesis testing and create 95% confidence intervals before exponentiating to transform back to the original scale.
Now consider the relative association:
$RR=\frac{\Psi(\mathbb{P};1)}{\Psi(\mathbb{P};0)}=\frac{\mu_{1}/\mu_{0}}{\mu_{3}/\mu_{2}}$ (4)
with corresponding subscripts for the influence curves of the TMLEs. Then the influence curve for the $log(RR)$ is given by
$IC_{RR}=(-1/\mu_{0})\times IC_{0}+(1/\mu_{1})\times IC_{1}+(1/\mu_{2})\times IC_{2}-(1/\mu_{3})\times IC_{3}$ (5)
We obtain inference as above.
Appendix S5: Simulation Study
We conducted a simulation study designed to mimic the causal model given in Figure 3 in the main text. The data generating process is as follows where $\text{expit}(x)=(1+e^{-x})^{-1}$.
  • Baseline covariates: $L_{0,1}\sim\text{Bernoulli}(0.5)$, $L_{0,2}\sim\text{Uniform}(0,1)$, and $L_{0,3}\sim\text{Uniform}(0,1)$.
  • Measurement indicator for the exposure: $\Delta_{A}\sim\text{Bernoulli}(\text{expit}(1.7+0.4L_{0,1}-0.5L_{0,3}))$
  • Exposure: $A\sim\text{Bernoulli}(\text{expit}(-0.3+1.5L_{0,1}-1.2L_{0,2}))$, observed only when $\Delta_{A}=1$.
  • Measurement indicator for the baseline outcome: $\Delta_{Y_{0}}\sim\text{Bernoulli}(\text{expit}(2.0+0.3L_{0,1}-0.4L_{0,2}-0.3A+0.2\Delta_{A}))$
  • Baseline outcome: $Y_{0}\sim\text{Bernoulli}(\text{expit}(-2.6+1.2L_{0,1}+0.8L_{0,2}-0.5L_{0,3}+0.7A+0.3A\cdot L_{0,1}))$, observed only when $\Delta_{Y_{0}}=1$.
  • Time-varying covariates: $L_{1,1}=\text{clip}(\mathcal{N}(0.3A+0.2L_{0,1}+0.15L_{0,2},\ 0.25^{2}),0,1)$ and $L_{1,2}\sim\text{Bernoulli}(\text{expit}(-0.3+1.2A+0.8L_{0,1}-0.6L_{0,3}+0.5Y_{0}))$
  • Measurement indicator for the endline outcome: $\Delta_{Y_{1}}\sim\text{Bernoulli}(\text{expit}(1.1+5.5L_{1,1}^{2}A-5.0L_{0,3}L_{1,1}L_{1,2}-4.0Y_{0}A+0.5L_{0,1}))$, among those with $\Delta_{Y_{0}}=1$ and equal to 0 otherwise.
  • Endline outcome: $Y_{1}\sim\text{Bernoulli}(\text{expit}(-2.0+5.0A\cdot L_{1,1}^{2}-4.5L_{0,3}^{2}L_{1,2}+1.5L_{0,1}+1.2A))$, observed only when $\Delta_{Y_{1}}=1$.
We calculated the true value of the exposure-specific estimands $\Psi(\mathbb{P};1)$ and $\Psi(\mathbb{P};0)$ as well as the relative and absolute effects by generating $10^{6}$ counterfactuals and taking the relevant means, ratios, and differences, respectively.
We implemented TMLE, G-computation, and inverse probability weighting (IPW) for the exposure-specific estimands and the effects. Mirroring the real-data analysis, we used Super Learner within TMLE to combine predictions from generalized linear models, the mean, and multivariate adaptive regression splines with and without covariate screening. We used parametric regressions in the singly robust approaches. Also mirroring the real-data analysis, we implemented a “Naïve-adjusted” approach: subset on participants with complete data ($\Delta_{A}=\Delta_{Y_{0}}=\Delta_{Y_{1}}=1$) and at-risk at baseline ($Y_{0}=0$) and then implemented TMLE with adjustment for confounding. For completeness in the simulations, we also implemented a “Naïve-unadjusted” approach: subset on participants with complete data ($\Delta_{A}=\Delta_{Y_{0}}=\Delta_{Y_{1}}=1$) and at-risk at baseline ($Y_{0}=0$) and then take the empirical mean outcome within exposure groups. Influence curve-based inference was used for variance estimation and confidence interval construction. We ran 1000 simulation studies at a sample size of $N=20000$ to reflect the motivating example.
Appendix S4.1: Simulation Results
The true values of the estimands were as follows: the expected outcome under the exposure $\Psi(\mathbb{P};1)=0.510$, the expected outcome under no exposure $\Psi(\mathbb{P};0)=0.177$, the risk ratio $RR=2.879$, and the risk difference $RD=0.333$. Table 1 provides results in terms of bias (average deviation between the point estimate and truth), empirical standard deviation (standard deviation [SD] of the point estimate across simulation replicates), the ratio of bias to empirical SD, and 95% confidence interval coverage (proportion of calculated intervals containing the truth). We report the bias-to-SD ratio as a diagnostic for the asymptotic behavior of our estimators: since the standard deviation shrinks at rate $\sqrt{N}$ while systematic bias remains fixed, a non-negligible bias-to-SD ratio indicates performance will deteriorate as sample size grows.
For all estimands, TMLE with Super Learner achieved nominal coverage, had negligible bias, and had bias-to-SD ratios well below 0.2. In contrast, G-computation and IPW only achieved nominal coverage for the expected outcome under the exposure $\Psi(\mathbb{P};1)$, but not the other estimands. Indeed, the coverage of G-computation was 93.2% for the expected outcome under no exposure $\Psi(\mathbb{P};0)$, 91.6% for the risk ratio, and 92.1% for the risk difference. The coverage of IPW was only 12.3% for the expected outcome under no exposure $\Psi(\mathbb{P};0)$, 19.7% for the risk ratio, and 58.9% for the risk difference.
Table 1: Over 1000 simulation iterations, estimator performance for TMLE, G-computation (“G-comp”), inverse probability weighting (IPW), and the naïve approach with and without confounding adjustment.
Estimand Estimator Bias Emp SD Bias/SD Coverage
$\Psi(\mathbb{P};1)$ TMLE 0.0014 0.0091 0.1560 0.9440
$\Psi(\mathbb{P};1)$ G-comp 0.0053 0.0083 0.6435 0.9500
$\Psi(\mathbb{P};1)$ IPW -0.0029 0.0089 -0.3231 0.9690
$\Psi(\mathbb{P};1)$ Naïve-adjusted 0.0565 0.0088 6.4347 0.0000
$\Psi(\mathbb{P};1)$ Naïve-unadjusted 0.1089 0.0080 13.5559 0.0000
$\Psi(\mathbb{P};0)$ TMLE 0.0003 0.0058 0.0578 0.9530
$\Psi(\mathbb{P};0)$ G-comp -0.0024 0.0054 -0.4546 0.9320
$\Psi(\mathbb{P};0)$ IPW 0.0179 0.0063 2.8376 0.1230
$\Psi(\mathbb{P};0)$ Naïve-adjusted 0.0102 0.0060 1.7023 0.6290
$\Psi(\mathbb{P};0)$ Naïve-unadjusted -0.0151 0.0052 -2.9191 0.0380
Risk Ratio TMLE 0.0056 0.1068 0.0525 0.9510
Risk Ratio G-comp 0.0733 0.1012 0.7247 0.9160
Risk Ratio IPW -0.2765 0.0953 -2.9025 0.1970
Risk Ratio Naïve-adjusted 0.1484 0.1054 1.4071 0.7240
Risk Ratio Naïve-unadjusted 0.9454 0.1316 7.1833 0.0000
Risk Difference TMLE 0.0011 0.0107 0.1012 0.9460
Risk Difference G-comp 0.0078 0.0097 0.7978 0.9210
Risk Difference IPW -0.0208 0.0109 -1.9127 0.5890
Risk Difference Naïve-adjusted 0.0463 0.0104 4.4549 0.0100
Risk Difference Naïve-unadjusted 0.1240 0.0095 13.0900 0.0000
We emphasize that TMLE, G-computation, and IPW rely on the same identification result. Therefore, differences in their performance are due to differences in statistical estimation. To explore further, Table 2 provides their performance for the numerator and denominator of our ratio-estimand:
$\Psi(\mathbb{P};a)=\frac{\mathbb{E}[\mathbb{E}\left(\mathbb{E}(Y_{1}\mid\Delta_{Y_{1}}=1,L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}\right)]}{1-\mathbb{E}\left[\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\right]}$
TMLE had negligible bias for all components; its bias-to-SD ratio remained below 0.2. In contrast, G-computation exhibited meaningful bias, as indicated by the bias-to-SD ratio, for all components except for the numerator under $A=1$. The numerator/denominator decomposition identifies the source IPW’s poor performance: the numerator for $A=0$ is substantially biased with a bias-to-SD ratio of $2.79$. This bias propagates directly into meaningful bias and poor coverage for the expected outcome under no exposure $\Psi(\mathbb{P};0)$, the risk ratio, and the risk difference.
Table 2: Further details on estimator performance with TMLE, G-comp, and IPW for the exposure-specific components of the ratio-estimands.
Estimator Exposure Component Truth Mean Est Bias Emp SD Bias/SD
TMLE $A=1$ Numerator 0.3593 0.3602 0.0009 0.0071 0.1274
TMLE $A=1$ Denominator 0.7047 0.7045 -0.0002 0.0056 -0.0346
TMLE $A=0$ Numerator 0.1499 0.1501 0.0002 0.0050 0.0419
TMLE $A=0$ Denominator 0.8464 0.8460 -0.0004 0.0044 -0.0957
G-comp $A=1$ Numerator 0.3593 0.3604 0.0011 0.0066 0.1686
G-comp $A=1$ Denominator 0.7047 0.6996 -0.0051 0.0054 -0.9529
G-comp $A=0$ Numerator 0.1499 0.1473 -0.0026 0.0046 -0.5549
G-comp $A=0$ Denominator 0.8464 0.8435 -0.0029 0.0044 -0.6663
IPW $A=1$ Numerator 0.3593 0.3571 -0.0021 0.0069 -0.3099
IPW $A=1$ Denominator 0.7047 0.7045 -0.0002 0.0056 -0.0344
IPW $A=0$ Numerator 0.1499 0.1649 0.0151 0.0054 2.7942
IPW $A=0$ Denominator 0.8464 0.8460 -0.0004 0.0045 -0.0950
Now we consider the naïve approaches, which are targeting a different (incorrect) estimand. Therefore, there was a substantial “causal gap”, even when adjusting for confounding.[53, 15, 14, 21, 48] By simply subsetting on participants with complete data and baseline risk, the naïve approaches failed to appropriately account for differential missingness. As shown in Table 1, the naïve approaches exhibited meaningful bias for all estimands. The coverage of naïve approach with confounding adjustment was 0% for the expected outcome under the exposure $\Psi(\mathbb{P};1)$, 62.9% for the expected outcome under no exposure $\Psi(\mathbb{P};0)$, 72.4% for the risk ratio, and 1% for the risk difference. Finally, the coverage of naïve approach without confounding adjustment was <<5% for all estimands.
Altogether, our simulation results support the robustness of TMLE with Super Learner in settings resembling our applied analysis. Our simulation results further suggest the practical differences between TMLE, G-computation, and IPW in SEARCH-TB are due to TMLE’s flexible adjustment for confounding and differential missingness. Again, all three approaches are targeting the same estimand. Therefore, the simulation and applied results highlight how theoretical properties (e.g., double robustness) translate into practical impact.
Appendix S6: Accounting for outcome dependence
Here, we outline an approach to account for the dependence of participant outcomes within groups or clusters, such as households, schools, hospitals, or communities. Such dependence could arise due to shared exposures and/or the spread of social behaviors or infectious diseases. In our running example, TB is transmitted from person to person. This dependence should be reflected in the corresponding causal model (e.g., [23, 24, 32, 6]). By following the Causal Roadmap or a similar framework for causal inference,[53, 27] we specify causal models encoding our knowledge about the hierarchical data generating process without imposing parametric modeling assumptions — in contrast to more traditional approaches, such as generalizing estimating equations or mixed effects models (e.g., [78, 6, 4, 50, 49]).
Suppose it is reasonable to assume that participant outcomes are dependent within households, but effectively independent between households. Then our causal model would be specified at the household-level, and identification would consider the influence of other household members as well as community-level factors. (See, for example, [6].) Concretely, this may involve including community indicators in $L_{0}$ and summary measures of household-level covariates in $L_{0}$ and $L_{1}$. The exact form of the causal model and identification result will depend on the application. Going forward, we use “cluster” to refer to any group considered to be the (conditionally) independent unit.[78, 50, 49]
If clustering is present, estimation and inference must be adjusted. First, the cross-validation scheme used within Super Learner must respect the independent unit. Concretely, participants in a given cluster are all assigned to the same sample-split. Second, variance estimation must account for clustering. Following Schnitzer et al.,[68] let $m=\{1,\ldots,M\}$ index the clusters and $j=\{1,\ldots,Z_{m}\}$ index for participants in cluster $m$. Then the total number of participants is $N=\sum_{m}Z_{m}$, and the asymptotic linearity result is re-expressed as $\hat{\Psi}-\Psi=\frac{1}{M}\sum_{m=1}^{M}\left(\sum_{j\in Z_{m}}D_{m,j}\frac{M}{N}\right)$ where $D_{m,j}$ denotes the influence curve for the $j^{th}$ participant in the $m^{th}$ cluster and where we suppressed the second-order remainder term for notational convenience.[68] Altogether, $X_{m}=\frac{M}{N}\sum_{j\in Z_{m}}D_{m,j}$ is the cluster-level influence curve, which has aggregated the individual-level influence curves within cluster $m$ and is weighted by the ratio of the number of clusters to the number of individuals $M/N$. We then proceed with variance estimation using the cluster-level influence curve. This approach is equivalent to using an independent working correlation matrix when obtaining robust (sandwich-based) inference. (As detailed [9], a slightly different formulation is needed if the causal estimand weights clusters, versus individuals, equally.)
Appendix S7: Missingness DAGs
There is a growing interest in the use of missingness graphs to represent studies with missingness on multiple variables and to assess whether causal effects can be identified (termed “recovered”).[43, 44] In particular, Moreno-Betancur and colleagues introduced complete-data DAGs (c-DAGs), missingness DAGs (m-DAGs), and their link.[46] They also provided a series of “canonical” m-DAGs for causal effects of point-treatment exposures.[46, 85] Holovchak et al. extended this work for the effects of longitudinal exposures, while questioning the practical utility of the m-DAG approach.[30] Specifically, they noted, “no general algorithms are available to decide on recoverability, and decisions have to be made on a case-by-case basis”.[30] As an alternative, we demonstrate how the Causal Roadmap, an established framework for causal inference, can used to define, identify, and estimate causal estimands, including the average treatment effect under missing exposures and Counterfactual Strata Effects.